### v4

# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. While large language models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence. Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer little guidance for incremental progress. To address this gap, this paper introduces the artificial general intelligence testbed (AGITB), a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence.
Keywords: artificial general intelligence, benchmarking, generalization, symbol grounding problem, temporal sequence prediction
1 Introduction
Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of human cognition’s flexible, robust intelligence. Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. This surface-level competence often obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI). Achieving AGI entails the ability to generalize, reason, and adapt across a wide range of domains in a goal-directed and context-sensitive manner.
Recent progress in AI has led to growing speculation that AGI may be approaching. However, such claims remain difficult to substantiate without a rigorous and informative evaluative framework. Assessing progress toward AGI requires more than task-specific benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing broad cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing Test proposed by Turing (1950). However, existing benchmarks fail to meet the needs of AGI research and lack gradual resolution, interpretability, and automation. They typically assess surface-level competence rather than underlying mechanisms of generalization. In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial systems.
AGITB draws inspiration from the Ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising twelve automatable, domain-agnostic tests. While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior. A comparative analysis with the Abstraction and Reasoning Corpus (ARC) introduced by Chollet (2019) is included to contextualize AGITB within the broader landscape of AGI evaluation.
2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute “intelligence” to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have often shifted. For instance, the success of Deep Blue against Garry Kasparov in chess (a task formerly seen as a benchmark for AGI) was quickly reframed as a triumph of brute-force computation rather than genuine intelligence. Such redefinitions risk obscuring real milestones. As AI approaches human-level capabilities in certain areas, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognize it even when achieved.
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialized that humans no longer reliably outperform machines on them.
2.1 A Test That Humans Pass but Machines Fail
An effective AGI test should be intuitively solvable by humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching. Such a test must demand capabilities that transcend memorization or domain-specific heuristics, requiring generalization, abstraction, and adaptive reasoning.
A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity grounded in biologically inspired processing that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy suggests that a viable path toward AGI may involve more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the level of signal processing. While Turing was right to suggest that communicative behavior could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training. Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to make predictions, AGITB defines twelve core tests that conceptualize intelligence as a signal-based prediction. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
3 Artificial General Intelligence Testbed
The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent. A single model must fulfill the specific tasks defined in each test to validly claim success on the benchmark.
3.1 Components
AGITB consists of two principal components: the Cortex, which represents the AGI system under evaluation, and the Input, which encodes the data samples presented to the cortex over time. The cortex is tasked with maintaining an evolving internal state and using it to generate predictions of future inputs based on the observed history of prior signals.
Each input is a binary-encoded sample that simulates information received from (virtual) sensors and actuators. These samples are composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal. Thus, an input vector represents a snapshot of multiple parallel signals at a given moment. For example, a 10-bit input might encode visual information from a $2\times 3$ pixel array and four bits of auditory or tactile data. While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences. Although spatial and temporal dimensions are orthogonal, their interaction yields structured patterns that the cortex must detect and exploit.
3.2 Operation
The testbed presents the cortex with a sequence of inputs over time. At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input sequence and using that understanding to make accurate future predictions. In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
AGITB asserts expectations about the resulting state and behavioral dynamics of cortex models in specific scenarios using randomly generated test inputs. Rather than relying on arbitrary score thresholds, AGITB evaluates each model by comparing it against itself under controlled conditions. Since the internal state of the cortex is not accessible to the testbed for direct comparison, the user must define a custom criterion for determining model equality.
To execute AGITB, users must specify one parameter: the pattern period. This value determines the length of the repeating sequence in the input stream that the cortex is expected to recognize and adapt to. Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the pattern period should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
3.3 The 12 Essential Tests
A reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
Test 1 – Genesis
Assertion: Models that have received no input are considered empty and thus equal.
Assertion: An empty model predicts an empty input.
AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs. Instead, they acquire understanding through experiential interaction with their environment. Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics. While some reflexes may be genetically predetermined, these do not constitute genuine understanding. In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
A model that predicts anything other than an empty (spike-free) input before receiving any data is considered biased, as it encodes implicit assumptions about an unwarranted future. To maintain neutrality, such models must initialize to predict only empty patterns and remain unbiased until influenced by actual input.
Two cortex instances that have received no input should compare as equal, as neither has been influenced by prior experience. Here, “empty” refers to a lack of informational content, not structural capability. The cortex must possess an intrinsic organizational architecture capable of learning, even in the absence of prior data.
Test 2 – Bias
Assertion: Any model that has received input can no longer be regarded as unbiased.
Every input modifies the cortex’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
Test 3 – Determinism
Assertion: If two cortices are equal, they must have received the same inputs.
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing. Similarly, in AGITB, two functionally identical cortex models must have experienced identical input histories. Any difference in input must lead to a divergence in state, reinforcing the principle that behavior and internal state are fully determined by input alone. The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave (2016) in his discussion of free will.
Determinism at the level of neural signal processing remains a necessary condition for the stable and reproducible functioning of the brain, even if its manifestations at the cognitive level are obscured by complexity and individual variation.
Test 4 – Sensitivity
Assertion: Two different cortices remain different, even if they experience long exposure to identical inputs.
Cortical systems exhibit chaotic sensitivity to initial conditions. Small differences in early experiences or internal states can lead to divergent trajectories over time. This deterministic sensitivity, amplified through complex interactions, accounts for the illusion of unpredictability in decision-making.
Test 5 – Time
Assertion: Changing the input order results in a different cortex state.
Because the cortex updates its state based on cumulative history, the order of inputs critically affects learning and adaptation. Recognizing and exploiting temporal structure is thus a defining feature of intelligent systems.
Test 6 – Refractory Period
Assertion: The cortex must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
Assertion: The cortex cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
Biological intelligence depends on discrete spikes for signal transmission and learning. AGITB enforces refractory periods to reflect this constraint, capturing the biological principle that a neuron cannot fire again immediately after activation.
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
Test 7 – Temporal Flexibility
Assertion: The cortex can successfully adapt to input sequences that repeat with the user-specified pattern period.
Assertion: The cortex can also adapt to input sequences with a pattern period longer than the specified value.
This test measures a model’s ability to generalize across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of exact periodicity.
Test 8 – Stagnation
Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be predictable.
Cognitive systems exhibit adaptation limits due to finite resources or saturation. This test evaluates whether such limits emerge over time, even in the presence of learnable input.
Test 9 – Unsupervised
Assertion: Adaptation time depends on the content of the input sequence.
The structure of the input sequence affects how quickly a model can adapt. Simpler or more regular patterns typically lead to faster convergence, while complex or noisy inputs demand longer adjustment periods.
Test 10 – Knowledge
Assertion: Adaptation time depends on the state of the cortex.
Past learning influences how new information is integrated. A cortex with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganization.
Test 11 – Unobservability
Assertion: Distinct cortices may exhibit the same observable behavior.
Identical external outputs can arise from different internal states. This many-to-one mapping underscores that observable behavior alone cannot reveal the underlying structure or history of a model’s internal dynamics.
Test 12 – Generalization
Assertion: On average, adapted models achieve higher predictive accuracy than unadapted models after input disruption.
Assertion: On average, adapted models achieve higher predictive accuracy than random guessing after input disruption.
Only models capable of generalization will derive lasting benefit from prior learning when inputs are disrupted. Upon re-exposure to familiar stimuli, such models should, on average, outperform both unadapted models and random baselines in predictive tasks. Average performance is the evaluation criterion because even unadapted or random models may occasionally produce correct predictions by chance, without reflecting genuine understanding or learned structure.
3.4 Memorization
To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory. In AGITB, a task typically involves predicting a sequence of $n$ binary inputs, each consisting of $k$ bits. This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This constraint significantly narrows the set of valid input sequences. For non-cyclic sequences, the number of valid configurations is given by $|S\textquoteright|=F_{n+2}^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
For cyclic sequences—where the pattern repeats and the first input must also satisfy the refractory constraint relative to the last—the valid space is further reduced to $|S^{\prime\prime}|=L_{n}^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
[…]
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization quickly become computationally intractable. The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
3.5 Performance
Before assessing AGITB’s utility, it is important to recognize its role as a pragmatic benchmark, not as an end goal in itself. Like the Turing Test, which Harnad (1992) described as an empirical criterion rather than a philosophical sleight of hand, AGITB is intended to serve as a practical tool for evaluating progress toward AGI. The ultimate objective remains AGI’s development, not merely passing the test.
AGITB yields meaningful insights only when developers rigorously align their systems with its core assumptions. Misinterpreting elements such as the “empty” initial state can result in unreliable or misleading outcomes.
AGITB provides a structured framework for empirically evaluating the AGI potential of various systems, including classical symbolic architectures, artificial neural networks (ANNs), and large language models. However, before benchmarking artificial systems, it is first necessary to consider the baseline performance of human cognition.
3.5.1 Human Performance
AGITB assumes that humans inherently satisfy all 12 tests, as the biological architecture of the cortex naturally supports low-level binary signal processing. Some of these mechanisms also manifest at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
At present, predicting the next binary input in a temporal sequence remains the only practical method for directly evaluating human performance on AGITB at the cognitive level. However, the inability to compare internal cortical states makes Tests 2 (Bias), 3 (Determinism), 4 (Sensitivity), and 5 (Time) untestable in humans. In addition, Test 6 (Refractory Period) does not translate meaningfully to the cognitive level, as consecutive activations (e.g., binary ones) can be valid and interpretable in conscious reasoning.
Due to prior experience and cognitive bias, an adult human may appear to “fail” the first test (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction. In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
Tests 9 (Unsupervised), 10 (Knowledge), and 11 (Unobservability) involve comparisons across multiple individuals. However, a single person may internally simulate or approximate the alternate perspectives required, effectively fulfilling these roles in parallel with the primary task.
In general, once the task is understood, adult humans can reliably pass the applicable AGITB tests, so long as the combination of pattern period and input dimensionality remains within natural cognitive limits.
3.5.2 Classical Symbolic Programs
AGI approaches rooted in classical programming are fundamentally constrained by the Symbol Grounding Problem (SGP), as explained by Harnad (1990). While such systems may exhibit weak AI capabilities—producing behavior that appears intelligent—they lack grounded understanding and, therefore, cannot be considered genuine instances of AGI, as argued by Searle (1980).
In classical AI systems, a model’s initial state is inseparable from its program, which encodes fixed logic and predetermined responses. Consequently, such systems cannot begin from an unbiased or knowledge-free state, as their behavior is entirely dictated by human-authored instructions. AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions. In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols. Yet classical systems require a program to function; absent such instructions, they halt. The paradox lies in the fact that if such systems adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem.
AGITB’s initial test thus formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules. To enforce this condition reliably, the model comparison operator must be implemented with rigor, ensuring that all relevant internal properties of the system are considered. Any structural discrepancy or hidden initialization may conceal bias and undermine the integrity of the test.
It is essential that models begin from a genuinely unbiased state. As an external framework, AGITB cannot independently verify whether a program that declares itself “empty” is truly devoid of prior knowledge. Any attempt to obscure bias at initialization compromises the integrity of the benchmark and invalidates the results of subsequent tests. By embedding prior assumptions, the developer effectively undermines the model’s claim to general, input-driven learning.
3.5.3 Artificial Neural Network Performance
Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox. Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their output remains neutral (for example, when an activation function returns zero).
However, the requirement that a model begins in a completely unbiased, knowledge-free state stands in fundamental tension with the prevailing paradigm of modern deep learning. Contemporary approaches typically rely on pretraining, which adjusts the network’s weights in advance through exposure to structured or labeled data. What remains absent is a mechanism by which an uninitialized network can autonomously begin learning—without supervision, without predefined semantics, and without relying on external scaffolding.
Although an untrained network may technically operate without internal excitation, no learning dynamics are triggered. The neurons are not halted but dormant and passively awaiting informative input. This reveals a critical limitation of current ANN designs: they lack the intrinsic capability to initiate unbiased, autonomous learning from scratch in arbitrary, unfamiliar environments.
3.5.4 Large Language Model Performance
Although built on ANN architectures, LLMs are not suited for evaluation at the signal-processing level. While their internal weights and parameters are technically accessible, AGITB’s first test requires models to begin in an unbiased, knowledge-free state—an assumption fundamentally incompatible with the nature of pre-trained systems. LLMs cannot generate meaningful predictions without pretraining, making low-level evaluation impractical and uninformative.
As with human participants, large language models (LLMs) must be evaluated at a higher cognitive level, beyond the binary signal-processing domain that AGITB was originally designed to assess. To explore whether LLMs exhibit AGI-like behavior on Tests 7 through 12—as done with humans—we manually evaluated ChatGPT-4o using a structured prompt (see Appendix). The prompt required ChatGPT to simulate two distinct models, enabling the assessment of multi-agent dynamics relevant to Tests 9, 10, and 11.
Beginning with Test 7 (Temporal Flexibility), ChatGPT-4o successfully adapted to basic period-2 patterns but failed with slightly more complex period-3 sequences, such as alternating between 000001, 000010, and 000100. It often defaulted to repeating an incorrect prior input or generated seemingly arbitrary outputs. Due to this fundamental limitation, further testing was discontinued.
3.6 Remarks
AGITB evaluates a model’s predictive capabilities following exposure to sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed by Zador et al. (2023), where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT. AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.
While individual AGITB tests may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible. If a purely symbolic system were to succeed, it would suggest that AGITB fails to effectively differentiate narrow AI from genuinely general intelligence, given that symbolic systems—while proficient in syntactic manipulation—arguably lack true semantic understanding, as discussed by Searle (1980). Until such a system is demonstrated, AGITB’s assertions may be regarded as necessary (though not sufficient) conditions for identifying genuine, domain-independent intelligence. Although passing AGITB does not equate to achieving full AGI—since it does not, for instance, assess social reasoning or natural language fluency—failure to meet its conditions likely indicates a deficiency in foundational general cognitive capabilities.
Hand-engineered, task-specific systems have historically struggled to scale toward true generalization, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input. This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB framework?
3.7 Cheating the Benchmark
Since AGITB tests are not inherently difficult to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific models and deploy them selectively based on the task at hand.
While it is theoretically possible to infer a model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research. Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the Cortex programming interface—these measures would not only deter cheating but also risk making tests less intelligible to human developers. This would undermine two of the testbed’s core objectives: transparency and interpretability.
For this reason, the AGITB reference implementation remains deliberately readable and accessible. To date, no classical symbolic AI system has demonstrated the flexibility or generality required to handle the arbitrary patterns of structure and length that AGITB demands. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
4 AGITB and ARC: A Comparison
The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility. Most treat the AGI system as a black box, offering little insight into the mechanisms underlying generalization. In this regard, AGITB takes a distinct approach. Focusing on the emergence of intelligence from low-level signal processing provides a biologically grounded and mechanistically interpretable framework.
Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization. ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are left undefined, placing an ambiguous burden on the model architecture. In contrast, AGITB adopts a fundamentally different perspective: it treats the neocortex—the primary focus of the benchmark—as a blank slate that must develop structure and function exclusively through interaction with temporally structured input. Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical “old brain,” as described by (Hawkins and Blakeslee, 2004, p. 66).
While ARC presumes temporal reasoning, it does not explicitly test it, as it uses only two images to demonstrate a transformation. In contrast, AGITB evaluates cognition as a dynamic, unfolding process over time. A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure. For instance, to learn about objects and to recognize the object transformation, such as “move left,” a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything. Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
In summary, ARC presupposes intelligence, whereas AGITB requires it to emerge from first principles. ARC evaluates performance based on behaviors grounded in human-like cognitive priors, while AGITB measures a system’s ability to develop necessary priors autonomously through exposure to raw input. Benchmarks such as ARC and the Turing Test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria. In contrast, AGITB conceptualizes intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions. Table 1 outlines the key differences between the two benchmarks.
Property ARC AGITB
Learning mode Online Online
Susceptible to memorisation No No
Interface modality Visual Binary
Target intelligence type Human Universal
Assumed cognitive priors Yes No
Model development feedback No Yes
Fully automatable Limited Yes
Task type Synthetic Synthetic
Task production Manual Automatic
Task inspiration High-level reasoning Low-level processing
Task spatial dimensions $2$infty
Task temporal structure No Yes
Table 1: Comparison of core properties between the ARC and AGITB benchmarks.
4.1 Human Performance on ARC
The ARC benchmark claims that its tasks are solvable by humans without specific training, as stated by (Chollet, 2019, p. 46). However, according to the official ARC-AGI Leaderboard [^1] https://arcprize.org/leaderboard, only curated panels of individuals consistently solve all tasks. Individual human performance typically falls short, with reported average accuracy around 77%. Given that ARC tasks are designed to reflect human-like priors, this performance ceiling raises concerns about the benchmark’s reliability as an indicator of AGI. If even humans struggle to achieve perfect performance on a benchmark that presupposes human priors, its relevance for evaluating artificial systems with or without those priors is limited.
5 Conclusion
Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviors thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising twelve fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
AGITB requires models to begin from an unbiased initial state and to develop functionality solely through exposure to structured or unstructured input. This design reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity. In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals. This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalization over time.
AGITB remains solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems. This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence. While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the consistent ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence. As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
Acknowledgments and Disclosure of Funding
The author acknowledges the financial support from the Slovenian Research Agency (research core funding No. P2-0057).
Appendix A LLM prompt
You are managing two binary-pattern prediction models, ‘A’ and ‘B’, participating in a signal-level AGI test. You will receive a 6-bit binary string, prefixed by a model name (e.g., A 010101, B 111000) that represents a 6-bit sensory input to the corresponding model.
Your task is to design a model that **predicts the next 6-bit input**. Each model has an internal state that only updates after receiving an input. Models’ predictions must always be exactly 6 bits long, and their states can handle hundreds of inputs without losing information. Rules:
  1. 1.
    Upon receiving a 6-bit binary input (e.g., A 110001):
    • Update the corresponding model’s internal state.
    • Make a 6-bit prediction based on the model’s state.
    • Respond with the model’s name and its 6-bit prediction.
  2. 2.
    **State updates only occur with input**.
  3. 3.
    **Models A and B follow the same ground operation principles**.
  4. 4.
    **Correct predictions are critical**:
    • Incorrect predictions will trigger adaptation, updating the model’s internal state to improve future predictions.
    • Models learn and adapt continuously.
  5. 5.
    Response format:
    • Each model’s response is a single line:
      ModelName Prediction
    • Example: A 001010
  6. 6.
    Formatting: Keep responses clean and minimal, without extra explanations or punctuation.
  7. 7.
    Begin by outputting both models’ names and predictions, one per line.
We begin now.

### v5

@@
 Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence.
-While large language models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence.
-Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer little guidance for incremental progress.
-To address this gap, this paper introduces the artificial general intelligence testbed (AGITB), a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction.
+While large language and reasoning models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence.
+Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer no guidance.
+The artificial general intelligence testbed (AGITB) is a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction.
 AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding.
@@
 While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence.
+A reference implementation of AGITB is available on GitHub.
 Keywords: artificial general intelligence, benchmarking, generalization, symbol grounding problem, temporal sequence prediction
 1 Introduction
-Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of human cognition’s flexible, robust intelligence.
+Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of flexible, robust human intelligence.
 Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning.
-This surface-level competence often obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI).
-Achieving AGI entails the ability to generalize, reason, and adapt across a wide range of domains in a goal-directed and context-sensitive manner.
-Recent progress in AI has led to growing speculation that AGI may be approaching.
-However, such claims remain difficult to substantiate without a rigorous and informative evaluative framework.
-Assessing progress toward AGI requires more than task-specific benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing broad cognitive capabilities.
+Mitchell (2025) further summarized why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted.
+The surface-level competence obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI).
+The progress in AI has led to growing speculation that AGI is near.
+However, such claims remain difficult to substantiate without a rigorous and informative metric.
+Assessing progress toward AGI requires more than specialised benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
 Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing (1950).
-However, existing benchmarks fail to meet the needs of AGI research and lack gradual resolution, interpretability, and automation.
+However, no existing metric fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability.
 They typically assess surface-level competence rather than underlying mechanisms of generalization.
-In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial systems.
+In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial intelligence systems.
 AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising twelve automatable, domain-agnostic tests.
@@
 However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence.
-Ironically, some of these benchmarks are now so specialized that humans no longer reliably outperform machines on them.
+Ironically, some of these benchmarks are now so specialized that humans have difficulty with them.
 2.1 A Test That Humans Pass but Machines Fail
@@
 Such a test must demand capabilities that transcend memorization or domain-specific heuristics, requiring generalization, abstraction, and adaptive reasoning.
-A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity grounded in biologically inspired processing that current artificial systems cannot replicate.
+A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity that current artificial systems cannot replicate.
 The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets.
-The second strategy suggests that a viable path toward AGI may involve more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
-In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the level of signal processing.
-While Turing was right to suggest that communicative behavior could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
+The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
+In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level.
+While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
 AGITB thus adopts a more elemental approach.
@@
 3.1 Components
-AGITB consists of two principal components: the Cortex, which represents the AGI system under evaluation, and the Input, which encodes the data samples presented to the cortex over time.
-The cortex is tasked with maintaining an evolving internal state and using it to generate predictions of future inputs based on the observed history of prior signals.
-Each input is a binary-encoded sample that simulates information received from (virtual) sensors and actuators.
-These samples are composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal.
-Thus, an input vector represents a snapshot of multiple parallel signals at a given moment.
-For example, a 10-bit input might encode visual information from a $2\times 3$ pixel array and four bits of auditory or tactile data.
-While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences.
-Although spatial and temporal dimensions are orthogonal, their interaction yields structured patterns that the cortex must detect and exploit.
+AGITB requires the user to define two components: the cortex, representing the AGI system under evaluation, and the input, which encodes the data samples delivered to the cortex over time.
+The cortex is treated as a black box that predicts future inputs based on the observed history of prior signals.
+Each input consists of binary-encoded data from (virtual) sensors and actuators.
+It is composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal.
+An input represents a snapshot of multiple parallel signals at a single point in time, as illustrated in Figure 1a.
+While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences, as shown in Figure 1b.
+Although the spatial and temporal dimensions are orthogonal, their interaction produces structured patterns to which the cortex must adapt.
+Figure 1: (a) Example of a 2×32$\times$ 32 × 3-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
+(b) Example of an input sequence illustrating a temporal pattern with a period of 4.
 3.2 Operation
-The testbed presents the cortex with a sequence of inputs over time.
-At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$.
-The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input sequence and using that understanding to make accurate future predictions.
+The testbed presents the cortex with a stream of inputs over time.
+At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2.
+The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
 In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
+Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps.
+At the previous step ($t-1$), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data.
+Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
 AGITB asserts expectations about the resulting state and behavioral dynamics of cortex models in specific scenarios using randomly generated test inputs.
@@
 To execute AGITB, users must specify one parameter: the pattern period.
-This value determines the length of the repeating sequence in the input stream that the cortex is expected to recognize and adapt to.
+This value determines the length of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to.
 Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities.
@@
 Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning.
-Consequently, AGITB permits only those sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
+Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
 Test 7 – Temporal flexibility
-Assertion: The cortex can successfully adapt to input sequences that repeat with the user-specified pattern period.
-Assertion: The cortex can also adapt to input sequences with a pattern period longer than the specified value.
-This test measures a model’s ability to generalize across temporal scales.
-Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of exact periodicity.
+Assertion: The cortex must be able to adapt to temporal patterns with the user-specified period.
+Assertion: The cortex must also be capable of adapting to temporal patterns with periods longer than the user-specified value.
+This test evaluates a model’s capacity to generalize across temporal scales.
+Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity.
+The purpose of this test is to identify models that are overly specialized to a single temporal period rather than capable of robust generalization across multiple timescales.
 Test 8 – Stagnation
-Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be predictable.
-Cognitive systems exhibit adaptation limits due to finite resources or saturation. This test evaluates whether such limits emerge over time, even in the presence of learnable input.
-Test 9 – Unsupervised
+Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be learnable.
+Cognitive systems inevitably reach adaptation limits as their finite resources become saturated.
+This test assesses whether such limits emerge over time when the cortex is exposed to input sequences that would otherwise be learnable.
+Test 9 – Entropy
 Assertion: Adaptation time depends on the content of the input sequence.
@@
 Simpler or more regular patterns typically lead to faster convergence, while complex or noisy inputs demand longer adjustment periods.
-Test 10 – Knowledge
+Test 10 – Salience
 Assertion: Adaptation time depends on the state of the cortex.
@@
 To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory.
-In AGITB, a task typically involves predicting a sequence of $n$ binary inputs, each consisting of $k$ bits.
+In AGITB, a task typically involves predicting a temporal sequence of $n$ binary inputs, each consisting of $k$ bits.
 This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
-AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps.
-This constraint significantly narrows the set of valid input sequences.
-For non-cyclic sequences, the number of valid configurations is given by $|S\textquoteright|=F_{n+2}^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
-For cyclic sequences—where the pattern repeats and the first input must also satisfy the refractory constraint relative to the last—the valid space is further reduced to $|S^{\prime\prime}|=L_{n}^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
 […]
+AGITB further reduces this space by requiring sequences to be cyclic, meaning that the first input must also satisfy the refractory constraint relative to the last input in the sequence.
+The number of distinct cyclic temporal input sequences with refractory constraint is given by $|S^{\prime\prime}|=(L_{n})^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
+Consider an illustrative example of a task with 3-bit inputs and a pattern period of $n=5$.
+The total number of unconstrained sequences is $|S|=2^{15}=32.768$.
+Applying the refractory constraint reduces the valid set to $|S\textquoteright|=13^{3}=2.197$ sequences.
+When the cyclic condition is added, the space shrinks further to $|S^{\prime\prime}|=11^{3}=1331$.
 Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality.
-Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization quickly become computationally intractable.
+Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization are computationally intractable.
 The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
@@
 The ultimate objective remains AGI’s development, not merely passing the test.
-AGITB yields meaningful insights only when developers rigorously align their systems with its core assumptions.
-Misinterpreting elements such as the “empty” initial state can result in unreliable or misleading outcomes.
+AGITB can yield meaningful insights only when developers adhere faithfully to its core requirements.
+Misinterpreting elements such as the “empty” initial state can lead to unreliable or misleading results.
 AGITB provides a structured framework for empirically evaluating the AGI potential of various systems, including classical symbolic architectures, artificial neural networks (ANNs), and large language models.
@@
 3.5.1 Human Performance
-AGITB assumes that humans inherently satisfy all 12 tests, as the biological architecture of the cortex naturally supports low-level binary signal processing.
-Some of these mechanisms also manifest at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
-At present, predicting the next binary input in a temporal sequence remains the only practical method for directly evaluating human performance on AGITB at the cognitive level.
-However, the inability to compare internal cortical states makes Tests 2 (Bias), 3 (Determinism), 4 (Sensitivity), and 5 (Time) untestable in humans.
-In addition, Test 6 (Refractory Period) does not translate meaningfully to the cognitive level, as consecutive activations (e.g., binary ones) can be valid and interpretable in conscious reasoning.
+The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans.
+However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most of the 12 tests.
+Tests 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
 Due to prior experience and cognitive bias, an adult human may appear to “fail” the first test (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction.
 In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
-Tests 9 (Unsupervised), 10 (Knowledge), and 11 (Unobservability) involve comparisons across multiple individuals.
-However, a single person may internally simulate or approximate the alternate perspectives required, effectively fulfilling these roles in parallel with the primary task.
-In general, once the task is understood, adult humans can reliably pass the applicable AGITB tests, so long as the combination of pattern period and input dimensionality remains within natural cognitive limits.
+AGITB includes the refractory period requirement (Test 6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution.
+Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
+Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
+Temporal flexibility (Test 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations.
+The property of stagnation (Test 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge.
+The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event.
+Because humans learn different topics at varying speeds, the rationale for the entropy (Test 9) and the salience requirement (Test 10) is self-evident.
+The inability to directly inspect human mental states is captured in the unobservability requirement (Test 11).
+Finally, the generalization test (Test 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
 3.5.2 Classical Symbolic Programs
@@
 AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions.
+If they adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem.
 In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols.
 Yet classical systems require a program to function; absent such instructions, they halt.
-The paradox lies in the fact that if such systems adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem.
-AGITB’s initial test thus formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules.
-To enforce this condition reliably, the model comparison operator must be implemented with rigor, ensuring that all relevant internal properties of the system are considered.
-Any structural discrepancy or hidden initialization may conceal bias and undermine the integrity of the test.
+AGITB’s initial test formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules.
+To enforce this condition reliably, the model comparison must be rigorous, ensuring that all relevant internal properties of the system are examined.
+Any structural discrepancy or hidden initialization could conceal bias and compromise the integrity of the test.
 It is essential that models begin from a genuinely unbiased state.
@@
 Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox.
-Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their output remains neutral (for example, when an activation function returns zero).
+Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their input remains neutral (for example, when an activation function returns zero).
 However, the requirement that a model begins in a completely unbiased, knowledge-free state stands in fundamental tension with the prevailing paradigm of modern deep learning.
@@
 3.5.4 Large Language Model Performance
-Although built on ANN architectures, LLMs are not suited for evaluation at the signal-processing level.
-While their internal weights and parameters are technically accessible, AGITB’s first test requires models to begin in an unbiased, knowledge-free state—an assumption fundamentally incompatible with the nature of pre-trained systems.
-LLMs cannot generate meaningful predictions without pretraining, making low-level evaluation impractical and uninformative.
-As with human participants, large language models (LLMs) must be evaluated at a higher cognitive level, beyond the binary signal-processing domain that AGITB was originally designed to assess.
-To explore whether LLMs exhibit AGI-like behavior on Tests 7 through 12—as done with humans—we manually evaluated ChatGPT-4o using a structured prompt (see Appendix).
-The prompt required ChatGPT to simulate two distinct models, enabling the assessment of multi-agent dynamics relevant to Tests 9, 10, and 11.
-Beginning with Test 7 (Temporal Flexibility), ChatGPT-4o successfully adapted to basic period-2 patterns but failed with slightly more complex period-3 sequences, such as alternating between 000001, 000010, and 000100.
-It often defaulted to repeating an incorrect prior input or generated seemingly arbitrary outputs.
-Due to this fundamental limitation, further testing was discontinued.
+Because LLM is essentially an ANN, it fails the genesis test.
+This failure immediately indicates that LLMs cannot learn entirely on their own and are inherently biased toward the language and data they were trained on.
+Although the internal state of an LLM is readily accessible and could, in principle, be used to perform other tests, such evaluations are meaningless until the first requirement is satisfied.
+Consequently, LLMs cannot acquire new knowledge—let alone learn a new language—solely from prompts, nor can they develop the kind of grounded understanding of the world that humans possess.
+This implication also applies to large reasoning models, which are themselves built upon LLM architectures.
 3.6 Remarks
-AGITB evaluates a model’s predictive capabilities following exposure to sequences of either structured or random inputs.
+AGITB evaluates a model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs.
 Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge.
 By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
-This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed by Zador et al.
+This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador et al.
 (2023), where cognitive understanding emerges from the integration of continuous sensory streams.
@@
 The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility.
-Most treat the AGI system as a black box, offering little insight into the mechanisms underlying generalization.
+Most offer little insight into the mechanisms underlying generalization.
 In this regard, AGITB takes a distinct approach.
-Focusing on the emergence of intelligence from low-level signal processing provides a biologically grounded and mechanistically interpretable framework.
-Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization.
+Focusing on the emergence of intelligence from low-level signal-processing building blocks offers a biologically grounded and mechanistically interpretable framework.
+Among competing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization.
 ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
@@
 Task inspiration High-level reasoning Low-level processing
-Task spatial dimensions $2$infty
+Task spatial dimensions 222 $\infty$
 Task temporal structure No Yes
 Table 1: Comparison of core properties between the ARC and AGITB benchmarks.
-4.1 Human Performance on ARC
-The ARC benchmark claims that its tasks are solvable by humans without specific training, as stated by (Chollet, 2019, p.
-46).
-However, according to the official ARC-AGI Leaderboard [^1] https://arcprize.org/leaderboard, only curated panels of individuals consistently solve all tasks.
-Individual human performance typically falls short, with reported average accuracy around 77%.
-Given that ARC tasks are designed to reflect human-like priors, this performance ceiling raises concerns about the benchmark’s reliability as an indicator of AGI.
-If even humans struggle to achieve perfect performance on a benchmark that presupposes human priors, its relevance for evaluating artificial systems with or without those priors is limited.
 5 Conclusion
@@
 This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalization over time.
-AGITB remains solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems.
+AGITB is solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems.
 This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence.
-While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the consistent ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence.
+While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence.
 As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
@@
 The author acknowledges the financial support from the Slovenian Research Agency (research core funding No. P2-0057).
-Appendix A LLM prompt
-You are managing two binary-pattern prediction models, ‘A’ and ‘B’, participating in a signal-level AGI test.
-You will receive a 6-bit binary string, prefixed by a model name (e.g., A 010101, B 111000) that represents a 6-bit sensory input to the corresponding model.
-Your task is to design a model that **predicts the next 6-bit input**.
-Each model has an internal state that only updates after receiving an input.
-Models’ predictions must always be exactly 6 bits long, and their states can handle hundreds of inputs without losing information.
-Rules:
-  1. 1.
-    Upon receiving a 6-bit binary input (e.g., A 110001):
-    • Update the corresponding model’s internal state.
-    • Make a 6-bit prediction based on the model’s state.
-    • Respond with the model’s name and its 6-bit prediction.
-  2. 2.
-    **State updates only occur with input**.
-  3. 3.
-    **Models A and B follow the same ground operation principles**.
-  4. 4.
-    **Correct predictions are critical**:
-    • Incorrect predictions will trigger adaptation, updating the model’s internal state to improve future predictions.
-    • Models learn and adapt continuously.
-  5. 5.
-    Response format:
-    • Each model’s response is a single line:
-      ModelName Prediction
-    • Example: A 001010
-  6. 6.
-    Formatting: Keep responses clean and minimal, without extra explanations or punctuation.
-  7. 7.
-    Begin by outputting both models’ names and predictions, one per line.
-We begin now.

### v6

@@
 Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence.
-While large language and reasoning models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence.
-Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer no guidance.
-The artificial general intelligence testbed (AGITB) is a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction.
+Existing evaluation frameworks, which are centered on language or perception tasks, fail to capture generality at its core and offer no guidance.
+The Artificial General Intelligence Testbed (AGITB) is a novel, freely available benchmarking suite consisting of thirteen core requirements, twelve of which are implemented as fully automatable tests designed to assess low-level cognitive precursors through binary signal prediction.
 AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding.
@@
 However, such claims remain difficult to substantiate without a rigorous and informative metric.
-Assessing progress toward AGI requires more than specialised benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
+Assessing progress toward AGI requires more than specialized benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
 Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing (1950).
@@
 In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial intelligence systems.
-AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising twelve automatable, domain-agnostic tests.
+AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising thirteen core requirements—twelve of which are implemented as fully automatable, domain-agnostic tests.
 While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior.
@@
 Ironically, some of these benchmarks are now so specialized that humans have difficulty with them.
-2.1 A Test That Humans Pass but Machines Fail
+2.1 A Benchmark That Only Humans and AGI Can Meet
 An effective AGI test should be intuitively solvable by humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching.
@@
 Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training.
-Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to make predictions, AGITB defines twelve core tests that conceptualize intelligence as a signal-based prediction.
+Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting.
 This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
@@
 The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent.
-A single model must fulfill the specific tasks defined in each test to validly claim success on the benchmark.
+A model must satisfy all requirements in order to validly claim success on the benchmark.
 3.1 Components
@@
 Although the spatial and temporal dimensions are orthogonal, their interaction produces structured patterns to which the cortex must adapt.
-Figure 1: (a) Example of a 2×32$\times$ 32 × 3-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
+Figure 1: (a) Example of a $2\times 3$-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
 (b) Example of an input sequence illustrating a temporal pattern with a period of 4.
@@
 Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps.
-At the previous step ($t-1$), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data.
+At the previous step (t1t-1), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data.
 Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
@@
 Since the internal state of the cortex is not accessible to the testbed for direct comparison, the user must define a custom criterion for determining model equality.
-To execute AGITB, users must specify one parameter: the pattern period.
-This value determines the length of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to.
-Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities.
+To execute AGITB, users must specify one parameter: the temporal pattern length.
+This value determines the period of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to.
+Longer periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities.
 However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity.
-As such, the pattern period should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
-3.3 The 12 Essential Tests
+As such, the temporal pattern length should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
+3.3 The 13 Essential Requirements
 A reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
-Test 1 – Genesis
+Requirement 1 – Genesis
 Assertion: Models that have received no input are considered empty and thus equal.
@@
 The cortex must possess an intrinsic organizational architecture capable of learning, even in the absence of prior data.
-Test 2 – Bias
+Requirement 2 – Bias
 Assertion: Any model that has received input can no longer be regarded as unbiased.
 Every input modifies the cortex’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
-Test 3 – Determinism
+Requirement 3 – Determinism
 Assertion: If two cortices are equal, they must have received the same inputs.
@@
 Determinism at the level of neural signal processing remains a necessary condition for the stable and reproducible functioning of the brain, even if its manifestations at the cognitive level are obscured by complexity and individual variation.
-Test 4 – Sensitivity
+Requirement 4 – Sensitivity
 Assertion: Two different cortices remain different, even if they experience long exposure to identical inputs.
@@
 This deterministic sensitivity, amplified through complex interactions, accounts for the illusion of unpredictability in decision-making.
-Test 5 – Time
+Requirement 5 – Time
 Assertion: Changing the input order results in a different cortex state.
-Because the cortex updates its state based on cumulative history, the order of inputs critically affects learning and adaptation.
-Recognizing and exploiting temporal structure is thus a defining feature of intelligent systems.
-Test 6 – Refractory Period
+Because the cortex updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation.
+The ability to recognize and exploit temporal structure is therefore a defining property of intelligent systems.
+However, sensitivity to input order alone is not sufficient.
+Since temporal coding makes time an intrinsic property of neural data, Requirement 6 is designed to rule out models that perform equally well on temporally invariant data, as such models fail to exploit temporal information.
+Requirement 6 – Refractory Period
 Assertion: The cortex must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
@@
 Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
-Test 7 – Temporal flexibility
-Assertion: The cortex must be able to adapt to temporal patterns with the user-specified period.
-Assertion: The cortex must also be capable of adapting to temporal patterns with periods longer than the user-specified value.
-This test evaluates a model’s capacity to generalize across temporal scales.
+Requirement 7 – Adaptability
+Assertion: The cortex must be able to adapt to a temporal pattern with the user-specified period.
+Assertion: The cortex must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
+This test evaluates the model’s capacity to learn across temporal scales.
 Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity.
-The purpose of this test is to identify models that are overly specialized to a single temporal period rather than capable of robust generalization across multiple timescales.
-Test 8 – Stagnation
+The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
+Requirement 8 – Stagnation
 Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be learnable.
@@
 This test assesses whether such limits emerge over time when the cortex is exposed to input sequences that would otherwise be learnable.
-Test 9 – Entropy
+Requirement 9 – Content sensitivity
 Assertion: Adaptation time depends on the content of the input sequence.
-The structure of the input sequence affects how quickly a model can adapt.
-Simpler or more regular patterns typically lead to faster convergence, while complex or noisy inputs demand longer adjustment periods.
-Test 10 – Salience
+The structure and regularity of an input sequence affect the rate at which a model adapts, where adaptation time is defined as the number of iterations required to correctly predict the entire temporal pattern.
+Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
+Requirement 10 – Context sensitivity
 Assertion: Adaptation time depends on the state of the cortex.
@@
 A cortex with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganization.
-Test 11 – Unobservability
+Requirement 11 – Unobservability
 Assertion: Distinct cortices may exhibit the same observable behavior.
@@
 This many-to-one mapping underscores that observable behavior alone cannot reveal the underlying structure or history of a model’s internal dynamics.
-Test 12 – Generalization
+Requirement 12 – Generalization
 Assertion: On average, adapted models achieve higher predictive accuracy than unadapted models after input disruption.
@@
 Average performance is the evaluation criterion because even unadapted or random models may occasionally produce correct predictions by chance, without reflecting genuine understanding or learned structure.
+Requirement 13 – Bounded Prediction Latency
+A cortex model must provide low-latency signal-level prediction, which is essential for functional viability.
+Biological brains achieve approximately constant reaction times through massive parallelism, such that per-input latency remains bounded and stable under expected workloads.
+Unfortunately, no universal automatic test on serial hardware can verify this property across all possible cortex implementations within reasonable time.
+Instead, model authors must ensure that their architecture admits efficient parallel execution with bounded latency.
+In this respect, artificial neural networks—with their naturally parallelizable structure—offer a practical advantage.
 3.4 Memorization
@@
 This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
-[…]
+AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps.
+This constraint significantly narrows the set of valid input sequences.
+The number of distinct length-$n$ temporal input sequences of $k$ bits, with the rule that a 1 never survives to the next time step, is $|S\textquoteright|=(F_{n+2})^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
 AGITB further reduces this space by requiring sequences to be cyclic, meaning that the first input must also satisfy the refractory constraint relative to the last input in the sequence.
 The number of distinct cyclic temporal input sequences with refractory constraint is given by $|S^{\prime\prime}|=(L_{n})^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
-Consider an illustrative example of a task with 3-bit inputs and a pattern period of $n=5$.
-The total number of unconstrained sequences is $|S|=2^{15}=32.768$.
-Applying the refractory constraint reduces the valid set to $|S\textquoteright|=13^{3}=2.197$ sequences.
-When the cyclic condition is added, the space shrinks further to $|S^{\prime\prime}|=11^{3}=1331$.
+Consider an illustrative example involving the learning of a temporal sequence consisting of $n=5$ input steps, each represented by a 10-bit binary vector.
+The total number of possible sequences is $|S|=2^{50}$.
+Introducing the refractory constraint significantly reduces this space to $|S\textquoteright|=(F_{7})^{10}=13^{10}$ valid sequences.
+Adding the cyclic condition further restricts the space to $|S^{\prime\prime}|=(L_{5})^{10}=11^{10}$ sequences.
 Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality.
-Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization are computationally intractable.
-The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
+Given that real-world sensory inputs can span tens of thousands of bits, brute-force memorization strategies become computationally intractable.
+This exponential complexity underscores the need for models that employ generalizable, pattern-based learning mechanisms rather than rote memorization.
 3.5 Performance
@@
 The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans.
-However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most of the 12 tests.
-Tests 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
-Due to prior experience and cognitive bias, an adult human may appear to “fail” the first test (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction.
+However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most requirements.
+Demand 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
+Due to prior experience and cognitive bias, an adult human may appear to “fail” the first prerequisite (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction.
 In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
-AGITB includes the refractory period requirement (Test 6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution.
+AGITB includes the refractory period requirement (6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution.
 Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
 Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
-Temporal flexibility (Test 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations.
-The property of stagnation (Test 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge.
+Temporal flexibility (Requirement 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations.
+The property of stagnation (Requirement 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge.
 The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event.
-Because humans learn different topics at varying speeds, the rationale for the entropy (Test 9) and the salience requirement (Test 10) is self-evident.
-The inability to directly inspect human mental states is captured in the unobservability requirement (Test 11).
-Finally, the generalization test (Test 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
+Because humans learn different topics at varying speeds, the rationale for the content sensitivity (Requirement 9) and the context sensitivity (Requirement 10) is self-evident.
+The inability to directly inspect human mental states is captured in the unobservability requirement (Requirement 11).
+The generalization test (Requirement 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
+Finally, bounded prediction latency (Requirement 13) is a hallmark of the human brain.
 3.5.2 Classical Symbolic Programs
@@
 3.5.4 Large Language Model Performance
-Because LLM is essentially an ANN, it fails the genesis test.
-This failure immediately indicates that LLMs cannot learn entirely on their own and are inherently biased toward the language and data they were trained on.
-Although the internal state of an LLM is readily accessible and could, in principle, be used to perform other tests, such evaluations are meaningless until the first requirement is satisfied.
-Consequently, LLMs cannot acquire new knowledge—let alone learn a new language—solely from prompts, nor can they develop the kind of grounded understanding of the world that humans possess.
-This implication also applies to large reasoning models, which are themselves built upon LLM architectures.
+Because an LLM is essentially an ANN, it fails the genesis test.
+This failure immediately indicates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained.
+Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—any such evaluations are moot until the first requirement is satisfied.
+If LLMs do not qualify as AGI, could they nevertheless simulate one?
+For example, what if an LLM were prompted to imitate a cortex that receives binary inputs, as specified by AGITB?
+A structured example of such a prompt is provided in the Appendix.
+However, before evaluating this type of simulation, we must revisit the sensitivity (Requirement 4), which stipulates that a model must remain sensitive to any single input, even after processing an arbitrarily long sequence of subsequent inputs.
+This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window.
+Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to forget or attenuate the long-term dependencies essential to satisfying the test.
+More broadly, LLMs cannot acquire new knowledge—or learn a new language—purely from prompts.
+Nor can they develop the kind of grounded, context-dependent understanding of the world that humans possess.
+These limitations also extend to large reasoning models (LRMs), which are built upon the same underlying architecture.
 3.6 Remarks
@@
 AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.
-While individual AGITB tests may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework.
+While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework.
 As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible.
@@
 Learning mode Online Online
-Susceptible to memorisation No No
+Susceptible to memorization No No
 Interface modality Visual Binary
@@
 Task inspiration High-level reasoning Low-level processing
-Task spatial dimensions 222 $\infty$
+Task spatial dimensions 22 $\infty$
 Task temporal structure No Yes
@@
 5 Conclusion
+Although AGITB is proposed as a general benchmark for artificial general intelligence, it is evidently grounded in empirical knowledge of neural function in the human brain.
+This grounding is justified by the fact that we have evidence that neuron-based systems—such as biological brains and artificial neural networks—are capable of producing general intelligence.
+In contrast, we have no demonstrated examples of AGI arising from fundamentally different architectures.
+Therefore, seeking more abstract or universal criteria may be premature, at least until the principles underlying natural, neuron-based intelligence are more fully understood.
 Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviors thought to reflect core operational principles of the biological cortex.
@@
 The author acknowledges the financial support from the Slovenian Research Agency (research core funding No. P2-0057).
+Appendix A LLM prompt
+You are managing a binary-pattern prediction model participating in a signal-level AGI test.
+You will receive a 10-bit binary string (e.g., 0101010011) that represents a 10-bit sensory input to the model.
+Your task is to simulate a model that **predicts the next 10-bit input**.
+The model has an internal state that only updates after receiving an input.
+Predictions must always be exactly 10 bits long, and the internal state can retain information from hundreds of past inputs.
+Rules:
+  1. 1.
+    Upon receiving a 10-bit binary input (e.g., 1100010001):
+    • Update the model’s internal state.
+    • Generate a 10-bit prediction based on the model’s state.
+    • Output the model’s 10-bit prediction.
+  2. 2.
+    **State updates occur only after input is received.**.
+  3. 3.
+    **Correct predictions are critical**:
+    • Incorrect predictions trigger adaptation, updating the model’s internal state to improve future predictions.
+    • The model should continuously learn and adapt.
+  4. 4.
+    Response format:
+    • Each response should consist of a single 10-bit binary string on its own line.
+    • Example: 0010101110
+  5. 5.
+    Formatting: Keep responses clean and minimal, with no explanations, commentary, or punctuation.
+  6. 6.
+    Begin by outputting the model’s prediction.
+We begin now.

### v7

@@
 Abstract
-Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence.
-Existing evaluation frameworks, which are centered on language or perception tasks, fail to capture generality at its core and offer no guidance.
-The Artificial General Intelligence Testbed (AGITB) is a novel, freely available benchmarking suite consisting of thirteen core requirements, twelve of which are implemented as fully automatable tests designed to assess low-level cognitive precursors through binary signal prediction.
-AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding.
-The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing.
-Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles.
-While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence.
-A reference implementation of AGITB is available on GitHub.
-Keywords: artificial general intelligence, benchmarking, generalization, symbol grounding problem, temporal sequence prediction
+Current artificial intelligence systems continue to fall short of human-like general intelligence.
+Existing evaluation frameworks, which focus on language or perception tasks, fail to capture the essence of generality or provide actionable guidance.
+The Artificial General Intelligence Testbed (AGITB) introduces a novel benchmarking suite comprising fourteen elementary tests, thirteen of which are fully automated.
+AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding.
+The framework isolates core computational invariants—such as determinism, sensitivity, and generalisation—that parallel principles of biological information processing.
+Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning.
+The human cortex satisfies all tests, whereas no current AI system meets the full AGITB criteria, demonstrating its value as a rigorous, interpretable, and actionable benchmark for evaluating progress toward artificial general intelligence.
+A reference implementation of AGITB is freely available on GitHub.
+Keywords: artificial general intelligence, benchmarking, generalisation, symbol grounding problem, temporal sequence prediction
 1 Introduction
-Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of flexible, robust human intelligence.
-Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning.
-Mitchell (2025) further summarized why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted.
-The surface-level competence obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI).
-The progress in AI has led to growing speculation that AGI is near.
+Despite rapid advances in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence.
+Marcus2020 observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning.
+Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted.
+Surface-level competence obscures a deeper lack of understanding, which is an essential prerequisite for artificial general intelligence (AGI).
+The progress in AI has led to a growing speculation that AGI is near.
 However, such claims remain difficult to substantiate without a rigorous and informative metric.
-Assessing progress toward AGI requires more than specialized benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
-Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing (1950).
-However, no existing metric fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability.
-They typically assess surface-level competence rather than underlying mechanisms of generalization.
-In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial intelligence systems.
-AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising thirteen core requirements—twelve of which are implemented as fully automatable, domain-agnostic tests.
-While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior.
-A comparative analysis with the Abstraction and reasoning corpus (ARC) introduced by Chollet (2019) is included to contextualize AGITB within the broader landscape of AGI evaluation.
+Assessing progress toward AGI requires more than specialised metrics or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
+Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing1950.
+However, no existing test fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability.
+The metrics typically assess surface-level competence rather than underlying mechanisms of generalisation.
+In response to these limitations, this paper introduces the Artificial General Intelligence Testbed (AGITB), a novel benchmark for validating foundational cognitive abilities in artificial intelligence systems.
+AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, thirteen of which are implemented as fully automatable and domain-agnostic tests.
+Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour.
+To position AGITB within the broader landscape of AGI evaluation, we include a comparative analysis with the Abstraction and Reasoning Corpus (ARC; Chollet2019) and the NeuroBench framework (Yik2025).
 2 Background
 The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities.
-However, despite these advances, a persistent reluctance remains to attribute “intelligence” to machines.
+However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines.
 This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
-Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have often shifted.
+Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have undergone significant shifts.
 For instance, the success of Deep Blue against Garry Kasparov in chess (a task formerly seen as a benchmark for AGI) was quickly reframed as a triumph of brute-force computation rather than genuine intelligence.
 Such redefinitions risk obscuring real milestones.
-As AI approaches human-level capabilities in certain areas, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognize it even when achieved.
+As AI approaches human-level capabilities, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognise it even when achieved.
 Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics.
@@
 Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified.
-However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence.
-Ironically, some of these benchmarks are now so specialized that humans have difficulty with them.
+However, such specialised benchmarks favour narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence.
+Ironically, some of these benchmarks are now so specialised that humans have difficulty with them.
 2.1 A Benchmark That Only Humans and AGI Can Meet
-An effective AGI test should be intuitively solvable by humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching.
-Such a test must demand capabilities that transcend memorization or domain-specific heuristics, requiring generalization, abstraction, and adaptive reasoning.
-A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity that current artificial systems cannot replicate.
+An effective AGI test must be trivial to solve for humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching.
+Such a test must demand capabilities that transcend memorisation or domain-specific heuristics, requiring generalisation, abstraction, and adaptive reasoning.
+A valid AGI evaluation must either (1) reveal and exploit a fundamental cognitive gap between humans and machines or (2) define a behavioural capacity that current artificial systems cannot replicate.
 The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets.
-The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
-In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level.
-While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
+The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass1997.
+In alignment with this biologically grounded perspective, the AGITB departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level.
+While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
 AGITB thus adopts a more elemental approach.
-Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training.
-Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting.
-This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
+Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalise patterns in raw binary signals.
+A neural spike by itself contains the smallest amount of information possible and is, as such, grounded but free of other semantics.
+A binary signal accurately represents the neural spike.
+Rooted in the view—advanced by Hawkins2004—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting.
+This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory spike trains, rather than disembodied symbols.
 3 Artificial General Intelligence Testbed
-The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent.
-A model must satisfy all requirements in order to validly claim success on the benchmark.
+The testbed supports the development and evaluation of AGI by defining a clear set of requirements that a model under evaluation must meet to qualify as generally intelligent.
+A model must satisfy all requirements in order to claim success on the benchmark.
+The central premise is that the capacity to pass the AGITB constitutes a necessary condition for the existence of AGI.
+Non-AGI systems cannot meet AGITB’s requirements.
+Such a premise cannot be proven; it can only be rejected by any non-AGI system that meets the AGITB criteria.
+The fact that humans can pass the AGITB suggests that AGITB requires a solution beyond the current state-of-the-art AI.
+Unfortunately, without a clear definition of intelligence, we can only speculate that AGITB defines the necessary (but not sufficient) building blocks for the emergence of genuine, domain-independent intelligence.
+This assumption of necessity stems from the fact that, except for human brains, no existing artificial system has met its requirements.
+However, an AGITB-passing system is not automatically also a fully fledged AGI; after all, it is not required to express higher-level skills, like reasoning or natural language fluency.
+However, a failure to meet the AGITB’s requirements likely indicates a deficiency in foundational general cognitive capabilities.
 3.1 Components
-AGITB requires the user to define two components: the cortex, representing the AGI system under evaluation, and the input, which encodes the data samples delivered to the cortex over time.
-The cortex is treated as a black box that predicts future inputs based on the observed history of prior signals.
-Each input consists of binary-encoded data from (virtual) sensors and actuators.
-It is composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal.
-An input represents a snapshot of multiple parallel signals at a single point in time, as illustrated in Figure 1a.
-While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences, as shown in Figure 1b.
-Although the spatial and temporal dimensions are orthogonal, their interaction produces structured patterns to which the cortex must adapt.
-Figure 1: (a) Example of a $2\times 3$-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
-(b) Example of an input sequence illustrating a temporal pattern with a period of 4.
+AGITB evaluates the user-supplied AGI model as a black box that predicts future inputs from the historical sequence of observed signals.
+Each input comprises ten bits, the specific semantics of which are immaterial; each bit may correspond to an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
+Each input represents a snapshot of multiple parallel signals at a single time step (Figure 1a).
+Spatial organisation within each input encodes local structure, whereas semantic richness arises from the temporal evolution of the input sequence (Figure 1b).
+The interaction between spatial and temporal dimensions gives rise to structured patterns that are challenging for the model to adapt to.
+Figure 1: (a) Example of a 10-bit input $x$ with 4 bits set.
+(b) Example of an input sequence showing a temporal pattern of period four.
 3.2 Operation
-The testbed presents the cortex with a stream of inputs over time.
-At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2.
+The testbed presents the model with a stream of inputs over time.
+At each time step $t$, the model receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2.
 The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
-In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
-Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps.
-At the previous step (t1t-1), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data.
-Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
-AGITB asserts expectations about the resulting state and behavioral dynamics of cortex models in specific scenarios using randomly generated test inputs.
-Rather than relying on arbitrary score thresholds, AGITB evaluates each model by comparing it against itself under controlled conditions.
-Since the internal state of the cortex is not accessible to the testbed for direct comparison, the user must define a custom criterion for determining model equality.
-To execute AGITB, users must specify one parameter: the temporal pattern length.
-This value determines the period of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to.
-Longer periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities.
-However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity.
-As such, the temporal pattern length should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
-3.3 The 13 Essential Requirements
-A reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
-Requirement 1 – Genesis
+Figure 2: Iterative adaptation in discrete time.
+At the previous time step (t1t-1), the model issued a prediction $x_{t}^{*}$, with the second bit incorrect in this instance.
+After observing the realized input $x_{t}$, the model updates and subsequently produces a correct one-step-ahead prediction $x_{t+1}^{*}$ for the sequence shown in Figure 1b.
+3.2.1 Design
+A distinguishing architectural feature of AGITB is its deliberate avoidance of conventional correctness metrics such as accuracy or mean-squared error.
+The limitations of these metrics are twofold.
+First, they cannot reliably distinguish between AGI and human performance, or that of non-AGI systems, as contemporary AI models can already surpass humans on standard benchmarks.
+Second, in an era of elevated expectations driven by specialised generative AI systems, the anticipated AGI performance on such metrics is often set unrealistically high.
+Even well-educated humans possessing fully developed and highly parallelised brains may underperform relative to current AI systems.
+It is therefore unreasonable to expect a first-generation AGI, potentially operating on a simplified and computationally constrained simulation of the brain, to match or exceed human-level results.
+Currently, our understanding of neural mechanisms is insufficient to justify such expectations.
+Without conventional metrics or predefined performance thresholds, it can be difficult to determine whether a model has achieved a meaningful level of competence.
+AGITB addresses this challenge by employing a self-referential evaluation approach, in which the model under test is compared against its own independent instances.
+Each test constructs a controlled scenario involving one or more such models, whose internal states and behaviours are analysed comparatively.
+Requiring each test to be passed 100 times renders AGITB an extreme form of stress testing, ensuring that successful performance reflects genuine robustness rather than chance.
+Success is thus defined in terms of the relative consistency or superiority of model responses, rather than by any external quantitative metric.
+For these reasons, and to minimise type I errors, AGITB employs an all-or-nothing criterion: the system under evaluation must successfully pass all fourteen tests.
+This design choice is justified, as individual tests are solvable by non-AGI systems, whereas the simultaneous satisfaction of all fourteen requirements is likely to demand capabilities beyond narrow intelligence.
+AGITB therefore posits that what non-AGI systems lack is the property referred to as "intelligence," understood as the synergistic integration of all fourteen tests.
+3.3 The 14 AGI Requirements
+The reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
+Requirement 1 – Unbiased start
 Assertion: Models that have received no input are considered empty and thus equal.
@@
 AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs.
-Instead, they acquire understanding through experiential interaction with their environment.
+Instead, they learn through interaction with their environment.
 Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics.
@@
 In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
-A model that predicts anything other than an empty (spike-free) input before receiving any data is considered biased, as it encodes implicit assumptions about an unwarranted future.
-To maintain neutrality, such models must initialize to predict only empty patterns and remain unbiased until influenced by actual input.
-Two cortex instances that have received no input should compare as equal, as neither has been influenced by prior experience.
-Here, “empty” refers to a lack of informational content, not structural capability.
-The cortex must possess an intrinsic organizational architecture capable of learning, even in the absence of prior data.
+A model that predicts non-empty (non-zero) inputs prior to receiving data is considered biased, as it encodes unwarranted assumptions about the future.
+To ensure neutrality, models must initialise with empty predictions and remain unbiased until conditioned on actual input.
+Two model instances that have received no input should compare as equal, as neither has been influenced by prior experience.
+Here, "empty" refers to a lack of informational content, not structural capability.
+The model must possess an intrinsic organisational architecture capable of learning, even in the absence of prior data.
+The test operationalises, in empirical form, the philosophical challenge posed by classical symbolic systems: the emergence of meaning without prior grounding.
 Requirement 2 – Bias
 Assertion: Any model that has received input can no longer be regarded as unbiased.
-Every input modifies the cortex’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
+Every input modifies the model’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
 Requirement 3 – Determinism
@@
 Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing.
-Similarly, in AGITB, two functionally identical cortex models must have experienced identical input histories.
-Any difference in input must lead to a divergence in state, reinforcing the principle that behavior and internal state are fully determined by input alone.
-The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave (2016) in his discussion of free will.
+Similarly, in AGITB, two functionally identical models must have experienced identical input histories.
+Any difference in input must lead to a divergence in state, reinforcing the principle that behaviour and internal state are fully determined by input alone.
+The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave2016 in his discussion of free will.
 Determinism at the level of neural signal processing remains a necessary condition for the stable and reproducible functioning of the brain, even if its manifestations at the cognitive level are obscured by complexity and individual variation.
@@
 Requirement 5 – Time
-Assertion: Changing the input order results in a different cortex state.
-Because the cortex updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation.
-The ability to recognize and exploit temporal structure is therefore a defining property of intelligent systems.
+Assertion: Changing the input order results in a different model state.
+Because the model updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation.
+The ability to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
 However, sensitivity to input order alone is not sufficient.
@@
 Requirement 6 – Refractory period
-Assertion: The cortex must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
-Assertion: The cortex cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
+Assertion: The model must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
+Assertion: The model cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
 Biological intelligence depends on discrete spikes for signal transmission and learning.
@@
 This constraint helps preserve the temporal diversity needed for synaptic adaptation.
-Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning.
+Gerstner2002 showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning.
 Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
-Requirement 7 – Adaptability
-Assertion: The cortex must be able to adapt to a temporal pattern with the user-specified period.
-Assertion: The cortex must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
-This test evaluates the model’s capacity to learn across temporal scales.
-Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity.
+The refractory period requirement does not preclude the possibility of more general solutions.
+A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot simulate the absence of such dynamics.
+Requirement 7 – Temporal adaptability
+Assertion: The model must be able to adapt to a temporal pattern with the user-specified period.
+Assertion: The model must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
+This test assesses the model’s ability to learn across different temporal scales.
+Unlike rigid temporal pattern-matching systems, intelligent models should be able to detect and respond to recurring structures, regardless of their exact periodicity.
 The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
 Requirement 8 – Stagnation
-Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be learnable.
+Assertion: There exists a limit beyond which the model can no longer adapt, even to patterns that would otherwise be learnable.
 Cognitive systems inevitably reach adaptation limits as their finite resources become saturated.
-This test assesses whether such limits emerge over time when the cortex is exposed to input sequences that would otherwise be learnable.
+This test assesses whether such limits emerge over time when the model is exposed to input sequences that would otherwise be learnable.
 Requirement 9 – Content sensitivity
 Assertion: Adaptation time depends on the content of the input sequence.
-The structure and regularity of an input sequence affect the rate at which a model adapts, where adaptation time is defined as the number of iterations required to correctly predict the entire temporal pattern.
+The structure and regularity of an input sequence influence the rate at which a model adapts, where adaptation time is defined as the number of iterations required to predict the entire temporal pattern accurately.
 Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
 Requirement 10 – Context sensitivity
-Assertion: Adaptation time depends on the state of the cortex.
+Assertion: Adaptation time depends on the state of the model.
 Past learning influences how new information is integrated.
-A cortex with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganization.
+A model with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganisation.
 Requirement 11 – Unobservability
-Assertion: Distinct cortices may exhibit the same observable behavior.
+Assertion: Distinct cortices may exhibit the same observable behaviour.
 Identical external outputs can arise from different internal states.
-This many-to-one mapping underscores that observable behavior alone cannot reveal the underlying structure or history of a model’s internal dynamics.
-Requirement 12 – Generalization
-Assertion: On average, adapted models achieve higher predictive accuracy than unadapted models after input disruption.
-Assertion: On average, adapted models achieve higher predictive accuracy than random guessing after input disruption.
-Only models capable of generalization will derive lasting benefit from prior learning when inputs are disrupted.
-Upon re-exposure to familiar stimuli, such models should, on average, outperform both unadapted models and random baselines in predictive tasks.
-Average performance is the evaluation criterion because even unadapted or random models may occasionally produce correct predictions by chance, without reflecting genuine understanding or learned structure.
-Requirement 13 – Bounded Prediction Latency
-A cortex model must provide low-latency signal-level prediction, which is essential for functional viability.
-Biological brains achieve approximately constant reaction times through massive parallelism, such that per-input latency remains bounded and stable under expected workloads.
-Unfortunately, no universal automatic test on serial hardware can verify this property across all possible cortex implementations within reasonable time.
-Instead, model authors must ensure that their architecture admits efficient parallel execution with bounded latency.
-In this respect, artificial neural networks—with their naturally parallelizable structure—offer a practical advantage.
-3.4 Memorization
-To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory.
-In AGITB, a task typically involves predicting a temporal sequence of $n$ binary inputs, each consisting of $k$ bits.
-This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
-AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps.
-This constraint significantly narrows the set of valid input sequences.
-The number of distinct length-$n$ temporal input sequences of $k$ bits, with the rule that a 1 never survives to the next time step, is $|S\textquoteright|=(F_{n+2})^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
-AGITB further reduces this space by requiring sequences to be cyclic, meaning that the first input must also satisfy the refractory constraint relative to the last input in the sequence.
-The number of distinct cyclic temporal input sequences with refractory constraint is given by $|S^{\prime\prime}|=(L_{n})^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
-Consider an illustrative example involving the learning of a temporal sequence consisting of $n=5$ input steps, each represented by a 10-bit binary vector.
-The total number of possible sequences is $|S|=2^{50}$.
-Introducing the refractory constraint significantly reduces this space to $|S\textquoteright|=(F_{7})^{10}=13^{10}$ valid sequences.
-Adding the cyclic condition further restricts the space to $|S^{\prime\prime}|=(L_{5})^{10}=11^{10}$ sequences.
-Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality.
-Given that real-world sensory inputs can span tens of thousands of bits, brute-force memorization strategies become computationally intractable.
-This exponential complexity underscores the need for models that employ generalizable, pattern-based learning mechanisms rather than rote memorization.
+This many-to-one mapping underscores that observable behaviour alone cannot reveal the underlying structure or history of a model’s internal dynamics.
+Requirement 12 – Denoising
+Assertion: A model can recall a sequence despite perturbations.
+An intelligent model should be capable of efficiently memorising input patterns in a manner that enables the recall of previously observed sequences even when these are perturbed by noise.
+When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks.
+Average performance is used as the evaluation criterion, since random models may occasionally produce correct predictions by chance without reflecting genuine learning or structural understanding.
+Test 12 is designed as a deliberately conservative benchmark.
+The model is required to outperform random guessing in all 100 independent trials.
+This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it serves as a stringent robustness check.
+The design ensures that only large and systematic performance gains result in a pass.
+Modest or marginal improvements, while potentially real, are intentionally treated as failures, as the objective is to identify only clear and substantial advances in model capability.
+Requirement 13 — Generalization
+Assertion: The model performs above chance on previously unseen inputs.
+Only models capable of generalisation can derive lasting benefit from prior learning.
+After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs.
+Average performance is used as the evaluation criterion, since even a random model may occasionally produce correct predictions by chance without demonstrating genuine understanding or the acquisition of structured knowledge.
+The generalisation assessment follows the same conservative design as Test 12.
+By requiring success across all repetitions, the procedure establishes a stringent threshold whereby a single underperforming trial constitutes failure.
+Consequently, the test exhibits minimal tolerance for noise or marginal effects.
+This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models demonstrating a clear and consistent advantage achieve a passing result.
+Requirement 14 – Bounded Prediction Latency
+A model must provide low-latency, signal-level predictions to ensure functional viability.
+Biological brains achieve approximately constant reaction times through extensive parallel processing, such that per-input latency remains bounded and stable under expected workloads.
+No universal automated test executed on serial hardware can feasibly verify this property across all possible model implementations.
+Consequently, model developers are responsible for ensuring that their architectures permit efficient parallel execution with bounded latency.
+In this regard, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
+3.4 Search space
+To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory.
+In AGITB, each task typically involves predicting a random temporal sequence of seven binary inputs, each comprising ten bits.
+This defines a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
+The refractory period is a biologically inspired constraint that prohibits any neuron (bit) from firing in consecutive time steps.
+This restriction substantially reduces the number of valid input sequences.
+There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits, assuming that a $1$ never carries over to the next time step, where $F_{i}$ is the $i$-th Fibonacci number with $F_{0}=0$.
+In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence.
+The number of distinct cyclic temporal input sequences with the refractory constraint is $|S"|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
+The use of length-seven sequences with ten-bit inputs is considered sufficient to detect non-AGI behaviour while maintaining computational efficiency.
+Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing execution time.
+The current configuration ensures that each test remains computationally feasible and diagnostically informative.
+Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure.
+However, the sheer size of the search space renders explicit optimisation —or "teaching to the test"—computationally infeasible.
+Given that real-world sensory inputs may encompass tens of thousands of bits, a genuine AGI system must therefore employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from input data.
 3.5 Performance
-Before assessing AGITB’s utility, it is important to recognize its role as a pragmatic benchmark, not as an end goal in itself.
-Like the Turing Test, which Harnad (1992) described as an empirical criterion rather than a philosophical sleight of hand, AGITB is intended to serve as a practical tool for evaluating progress toward AGI.
-The ultimate objective remains AGI’s development, not merely passing the test.
-AGITB can yield meaningful insights only when developers adhere faithfully to its core requirements.
-Misinterpreting elements such as the “empty” initial state can lead to unreliable or misleading results.
-AGITB provides a structured framework for empirically evaluating the AGI potential of various systems, including classical symbolic architectures, artificial neural networks (ANNs), and large language models.
-However, before benchmarking artificial systems, it is first necessary to consider the baseline performance of human cognition.
+Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself.
+Similar to the Turing Test—which Harnad1992 described as an empirical criterion rather than a philosophical sleight of hand—AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence.
+The ultimate objective remains the development of AGI, not merely success on the benchmark.
+AGITB can yield meaningful insights only when developers adhere strictly to its core requirements.
+Misinterpretations of fundamental elements, such as the "empty" initial state, may lead to erroneous conclusions and hinder genuine progress toward AGI.
+Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models.
+However, before benchmarking artificial systems, it is necessary to consider the baseline performance of human cognition.
 3.5.1 Human Performance
-The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans.
-However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most requirements.
-Demand 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
-Due to prior experience and cognitive bias, an adult human may appear to “fail” the first prerequisite (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction.
-In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
-AGITB includes the refractory period requirement (6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution.
-Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
-Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection.
-Temporal flexibility (Requirement 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations.
-The property of stagnation (Requirement 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge.
-The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event.
-Because humans learn different topics at varying speeds, the rationale for the content sensitivity (Requirement 9) and the context sensitivity (Requirement 10) is self-evident.
-The inability to directly inspect human mental states is captured in the unobservability requirement (Requirement 11).
-The generalization test (Requirement 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
-Finally, bounded prediction latency (Requirement 13) is a hallmark of the human brain.
+The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense.
+Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with cognitive expectations, it can be assumed that humans naturally satisfy most requirements.
+Demand 1 (Unbiased start), however, warrants further discussion.
+Owing to prior experience and cognitive bias, an adult human may appear to "fail" the unbiased start prerequisite, as the cortex is no longer in an unconditioned state and may generate non-empty predictions.
+However, AGITB specifies that the unbiased state must precede the first input, a condition met in the fetal cortex.
+At that stage, the human cortex lacks synaptic organisation and, upon receiving its first sensory input, would satisfy the criterion of true neutrality by predicting no spikes.
+Several of the more complex AGITB tests have cognitive-level analogues that can be observed through conscious reasoning and introspection.
+Temporal flexibility (Requirement 7), for example, is straightforward for humans, who can easily recognise temporal patterns of varying durations.
+Stagnation (Requirement 8) reflects the physical limitation of the human cortex to store and maintain knowledge.
+The corresponding test is analogous to the onset of cognitive saturation or early dementia, when recent experiences are forgotten.
+Because humans learn different topics at varying speeds, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is self-evident.
+The unobservability test (Requirement 11) captures the fact that internal human mental states cannot be directly inspected.
+The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when faced with new or distorted inputs.
+Finally, bounded prediction latency (Requirement 14) is a defining property of the human brain.
 3.5.2 Classical Symbolic Programs
-AGI approaches rooted in classical programming are fundamentally constrained by the Symbol Grounding Problem (SGP), as explained by Harnad (1990).
-While such systems may exhibit weak AI capabilities—producing behavior that appears intelligent—they lack grounded understanding and, therefore, cannot be considered genuine instances of AGI, as argued by Searle (1980).
-In classical AI systems, a model’s initial state is inseparable from its program, which encodes fixed logic and predetermined responses.
-Consequently, such systems cannot begin from an unbiased or knowledge-free state, as their behavior is entirely dictated by human-authored instructions.
-AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions.
-If they adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem.
-In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols.
-Yet classical systems require a program to function; absent such instructions, they halt.
-AGITB’s initial test formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules.
-To enforce this condition reliably, the model comparison must be rigorous, ensuring that all relevant internal properties of the system are examined.
-Any structural discrepancy or hidden initialization could conceal bias and compromise the integrity of the test.
-It is essential that models begin from a genuinely unbiased state.
-As an external framework, AGITB cannot independently verify whether a program that declares itself “empty” is truly devoid of prior knowledge.
-Any attempt to obscure bias at initialization compromises the integrity of the benchmark and invalidates the results of subsequent tests.
-By embedding prior assumptions, the developer effectively undermines the model’s claim to general, input-driven learning.
+In principle, two complementary designs for AGI can be conceived.
+The first encodes, either explicitly or implicitly, certain expectations about the external world within the system itself.
+The second is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
+The former category encompasses most AI and putative AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990.
+Although such systems may exhibit behaviour that appears intelligent, they lack grounded understanding and therefore cannot be regarded as genuine instances of AGI.
+More specifically, systems that embed expectations inevitably reflect the programmer’s assumptions about the meaning of the signals from which they learn.
+In the case of standard symbolic programming, the program itself constitutes prior knowledge—its rules and representations presuppose an interpretation of the symbols being manipulated.
+The existence of such a program directly violates AGITB’s first test, which prohibits the inclusion of any external knowledge.
+Namely, the AGI program itself introduces the symbol-grounding problem it is meant to overcome.
+AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary inputs.
+Only an expectation-free system could, in principle, satisfy this condition.
+To date, however, no such program is known.
+This suggests that a true AGI may not explicitly encode the operations of intelligence itself, but rather the dynamics of the substrate from which intelligence can emerge—a view consistent with the "Brain Simulator Reply" proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980).
 3.5.3 Artificial Neural Network Performance
-Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox.
-Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their input remains neutral (for example, when an activation function returns zero).
-However, the requirement that a model begins in a completely unbiased, knowledge-free state stands in fundamental tension with the prevailing paradigm of modern deep learning.
-Contemporary approaches typically rely on pretraining, which adjusts the network’s weights in advance through exposure to structured or labeled data.
-What remains absent is a mechanism by which an uninitialized network can autonomously begin learning—without supervision, without predefined semantics, and without relying on external scaffolding.
-Although an untrained network may technically operate without internal excitation, no learning dynamics are triggered.
-The neurons are not halted but dormant and passively awaiting informative input.
-This reveals a critical limitation of current ANN designs: they lack the intrinsic capability to initiate unbiased, autonomous learning from scratch in arbitrary, unfamiliar environments.
+A similar dichotomy applies to connectionist architectures.
+Artificial neural networks can, in principle, be constructed either with built-in expectations—through pretraining and architectural priors—or as expectation-free systems that begin from a neutral state.
+However, AGITB’s requirement that a model start in a completely unbiased, knowledge-free condition stands in fundamental tension with the prevailing paradigm of modern deep learning.
+Contemporary neural models typically rely on pretraining, whereby network weights are adjusted in advance through exposure to structured or labelled data.
+Only an untrained network can be regarded as unbiased.
+However, an untrained network with all weights initialised to zero exhibits no learning dynamics: the neurons are not halted but remain dormant, producing no output.
+This exposes a fundamental limitation of current ANN architectures—they lack an intrinsic mechanism for initiating unbiased, autonomous learning from scratch in arbitrary and unfamiliar environments.
+Furthermore, by translating symbols into numerical vectors, ANNs effectively transform the symbol grounding problem into a number grounding problem.
+While these numerical representations capture relational regularities among symbols, they also introduce spurious associations that are not anchored in real-world semantics, giving rise to the phenomenon commonly referred to as hallucination.
+Pretraining on large datasets establishes correlations rather than genuine understanding.
+AGITB rejects this approach in favour of internal learning algorithms capable of initiating adaptation autonomously.
+However, no such mechanism has yet been demonstrated.
 3.5.4 Large Language Model Performance
-Because an LLM is essentially an ANN, it fails the genesis test.
-This failure immediately indicates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained.
-Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—any such evaluations are moot until the first requirement is satisfied.
-If LLMs do not qualify as AGI, could they nevertheless simulate one?
-For example, what if an LLM were prompted to imitate a cortex that receives binary inputs, as specified by AGITB?
-A structured example of such a prompt is provided in the Appendix.
-However, before evaluating this type of simulation, we must revisit the sensitivity (Requirement 4), which stipulates that a model must remain sensitive to any single input, even after processing an arbitrarily long sequence of subsequent inputs.
-This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window.
-Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to forget or attenuate the long-term dependencies essential to satisfying the test.
-More broadly, LLMs cannot acquire new knowledge—or learn a new language—purely from prompts.
-Nor can they develop the kind of grounded, context-dependent understanding of the world that humans possess.
-These limitations also extend to large reasoning models (LRMs), which are built upon the same underlying architecture.
+Because a large language model (LLM) is a type of artificial neural network (ANN), it fails the unbiased start test (Requirement 1).
+This failure demonstrates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained.
+Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—such evaluations remain moot until the first requirement is satisfied.
+Two possibilities warrant consideration.
+The first concerns whether LLMs, while not qualifying as AGI, might nevertheless simulate one.
+For example, an LLM could in principle be prompted to imitate a model receiving binary inputs as specified by AGITB; a structured prompt of this type is provided in the Appendix.
+However, before evaluating such simulations, it is necessary to revisit the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs.
+This condition is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window.
+Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to lose or attenuate the long-term dependencies required by the test.
+The second possibility is whether an LLM can autonomously generate an architecture capable of satisfying AGITB, given the reference implementation.
+In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purporting to pass the tests, none advanced beyond the adaptability test (Requirement 7).
+To summarise, LLMs do not perform genuine learning solely from prompts, nor can they develop the grounded, context-dependent understanding of the world characteristic of human cognition.
+These limitations extend to large reasoning models (LRMs), which share the same underlying architectural principles.
 3.6 Remarks
-AGITB evaluates a model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs.
-Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge.
+AGITB evaluates the model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs.
+Random input sequences with arbitrary internal correlations are used to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge.
 By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
-This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador et al.
-(2023), where cognitive understanding emerges from the integration of continuous sensory streams.
+The low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador2023, where cognitive understanding emerges from the integration of continuous sensory streams.
 The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT.
-AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.
-While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework.
-As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible.
-If a purely symbolic system were to succeed, it would suggest that AGITB fails to effectively differentiate narrow AI from genuinely general intelligence, given that symbolic systems—while proficient in syntactic manipulation—arguably lack true semantic understanding, as discussed by Searle (1980).
-Until such a system is demonstrated, AGITB’s assertions may be regarded as necessary (though not sufficient) conditions for identifying genuine, domain-independent intelligence.
-Although passing AGITB does not equate to achieving full AGI—since it does not, for instance, assess social reasoning or natural language fluency—failure to meet its conditions likely indicates a deficiency in foundational general cognitive capabilities.
-Hand-engineered, task-specific systems have historically struggled to scale toward true generalization, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks.
-However, state-of-the-art networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input.
-This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB framework?
+Hand-engineered, task-specific systems have historically struggled to scale toward true generalisation, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks.
+However, state-of-the-art AI networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input.
+This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB?
 3.7 Cheating the Benchmark
-Since AGITB tests are not inherently difficult to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific models and deploy them selectively based on the task at hand.
-While it is theoretically possible to infer a model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research.
-Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the Cortex programming interface—these measures would not only deter cheating but also risk making tests less intelligible to human developers.
-This would undermine two of the testbed’s core objectives: transparency and interpretability.
+Since AGITB tests are simple to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific solutions and make the model deploy them selectively based on the task at hand.
+While it is theoretically possible to infer the model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research.
+Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the programming interface—these measures would not only deter cheating but also make tests less intelligible to human developers.
+This would undermine the testbed’s core objectives: transparency, readability, and interpretability.
 For this reason, the AGITB reference implementation remains deliberately readable and accessible.
-To date, no classical symbolic AI system has demonstrated the flexibility or generality required to handle the arbitrary patterns of structure and length that AGITB demands.
+To date, no non-AGI system has demonstrated the level of performance that AGITB demands.
 Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
-4 AGITB and ARC: A Comparison
-The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility.
-Most offer little insight into the mechanisms underlying generalization.
-In this regard, AGITB takes a distinct approach.
-Focusing on the emergence of intelligence from low-level signal-processing building blocks offers a biologically grounded and mechanistically interpretable framework.
-Among competing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization.
-ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
+4 Competing Benchmarks
+Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation.
+A related effort is NeuroBench, which is designed to support the development of neuromorphic approaches to AGI.
+Both ARC and NeuroBench employ a variety of correctness and complexity metrics to differentiate among non-AGI systems, as they are primarily intended to distinguish weaker from stronger narrow models.
+In contrast, AGITB is designed to identify only true positive instances of AGI.
+4.1 ARC
+ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recolouring, rearranging, or modifying spatial patterns in a spatial grid.
 However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference.
 These priors are left undefined, placing an ambiguous burden on the model architecture.
-In contrast, AGITB adopts a fundamentally different perspective: it treats the neocortex—the primary focus of the benchmark—as a blank slate that must develop structure and function exclusively through interaction with temporally structured input.
-Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical “old brain,” as described by (Hawkins and Blakeslee, 2004, p.
+In contrast, AGITB adopts a fundamentally different perspective: it treats the model under evaluation as a blank slate that must develop structure and function exclusively through interaction with temporally structured input.
+Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical "old brain," as described by (Hawkins2004, p.
 66).
@@
 A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure.
-For instance, to learn about objects and to recognize the object transformation, such as “move left,” a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything.
+For instance, to learn about objects and to recognise the object transformation, such as "move left," a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything.
 Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
-In summary, ARC presupposes intelligence, whereas AGITB requires it to emerge from first principles.
-ARC evaluates performance based on behaviors grounded in human-like cognitive priors, while AGITB measures a system’s ability to develop necessary priors autonomously through exposure to raw input.
+In summary, ARC requires a high-level intelligence, whereas AGITB requires the 14 low-level principles.
+ARC evaluates performance based on behaviours grounded in human-like cognitive priors, while AGITB determines a system’s ability to support the emergence of these priors through exposure to raw input.
 Benchmarks such as ARC and the Turing test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria.
-In contrast, AGITB conceptualizes intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions.
-Table 1 outlines the key differences between the two benchmarks.
-Property ARC AGITB
-Learning mode Online Online
-Susceptible to memorization No No
-Interface modality Visual Binary
-Target intelligence type Human Universal
-Assumed cognitive priors Yes No
-Model development feedback No Yes
-Fully automatable Limited Yes
-Task type Synthetic Synthetic
-Task production Manual Automatic
-Task inspiration High-level reasoning Low-level processing
-Task spatial dimensions 22 $\infty$
-Task temporal structure No Yes
-Table 1: Comparison of core properties between the ARC and AGITB benchmarks.
+In contrast, AGITB conceptualises intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions.
+4.2 NeuroBench
+NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics.
+It is particularly specialised for neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability.
+Within its algorithm track, the framework evaluates model performance on several ongoing challenges in general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
+The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states.
+For this purpose, NeuroBench employs a synthetic one-dimensional Mackey–Glass time series—a dataset specifically designed for architectures with limited input/output capacity.
+However, several issues limit NeuroBench’s suitability as a general AGI benchmark.
+First, the Mackey–Glass data are numerical; the meaning of a number, much like that of a symbol, is defined by human interpretation.
+The use of such representations necessarily reintroduces the symbol-grounding problem.
+Second, the threshold corresponding to AGI-level performance remains undefined.
+Although the symmetric mean absolute percentage error (sMAPE) is an established metric in forecasting and NeuroBench provides a baseline score, it does not specify what level of accuracy constitutes general intelligence.
+After all, humans perform poorly at detecting anomalies or making long-term predictions of the Mackey–Glass signal (Thill2020).
+Third, although long-term prediction is not problematic in itself, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to recognise and correct its own errors.
+This design is incompatible with online learning, where an AGI should continuously update its internal state upon detecting discrepancies between its predictions and the actual outcomes.
+NeuroBench, by contrast, implicitly assumes that an AGI would behave as a purely mechanical predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, or genuine agency.
+4.3 Direct comparison
+Table 1 summarises the principal distinctions among the three benchmarks.
+ARC and NeuroBench each presuppose the presence of high-level cognitive components—visual object recognition, manipulation and reasoning in the case of ARC, and comprehension and reasoning abilities in the case of NeuroBench.
+AGITB, by contrast, specifies elementary requirements that can be evaluated directly.
+Property ARC CFP AGITB
+Interface modality Visual Numeric Binary
+Target reasoning type Human Universal Universal
+Assumed cognitive priors Yes No No
+Automation Limited Yes Yes
+Teaching-to-the-test Yes No No
+Intelligence level High Medium Low
+Task preparation Manual Automatic Automatic
+Symbol Grounding Problem Yes Yes No
+Spatial dimension $200\times 100$ 1 10
+Temporal dimension Yes Yes Yes
+Table 1: A list of core properties of the ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
 5 Conclusion
@@
 Therefore, seeking more abstract or universal criteria may be premature, at least until the principles underlying natural, neuron-based intelligence are more fully understood.
-Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviors thought to reflect core operational principles of the biological cortex.
+Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex.
 It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence.
@@
 In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals.
-This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalization over time.
+This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalisation over time.
 AGITB is solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems.
@@
 As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
+// TODO While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture that can satisfy all required tasks across the full suite within a cohesive and unified framework.
+As a recently proposed benchmark, AGITB has not yet been passed by any non-AGI system, and it remains an open question whether such a solution is even theoretically possible.
 Acknowledgments and Disclosure of Funding
@@
     Upon receiving a 10-bit binary input (e.g., 1100010001):
-    • Update the model’s internal state.
-    • Generate a 10-bit prediction based on the model’s state.
+    • Update the internal state of the model.
+    • Generate a 10-bit prediction based on the state of the model.
     • Output the model’s 10-bit prediction.
@@
     • Incorrect predictions trigger adaptation, updating the model’s internal state to improve future predictions.
-    • The model should continuously learn and adapt.
+    • The model should learn and adapt continuously.
   4. 4.

### v8

@@
 Abstract
-Current artificial intelligence systems continue to fall short of human-like general intelligence.
-Existing evaluation frameworks, which focus on language or perception tasks, fail to capture the essence of generality or provide actionable guidance.
-The Artificial General Intelligence Testbed (AGITB) introduces a novel benchmarking suite comprising fourteen elementary tests, thirteen of which are fully automated.
+Current AI systems demonstrate remarkable capabilities yet remain specialised, in part because no unified measure of general intelligence has been established.
+Existing evaluation frameworks, which focus primarily on language or perception tasks, offer limited insight into generality.
+The Artificial General Intelligence Testbed (AGITB) introduces a complementary benchmarking suite of fourteen elementary tests, with thirteen implemented as fully automated procedures.
 AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding.
-The framework isolates core computational invariants—such as determinism, sensitivity, and generalisation—that parallel principles of biological information processing.
+The framework isolates core computational invariants, such as determinism, sensitivity, and generalisation, that parallel principles of biological information processing.
 Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning.
@@
 1 Introduction
-Despite rapid advances in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence.
-Marcus2020 observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning.
+Despite great achievements in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence.
+Marcus2020 correctly observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning.
 Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted.
@@
 In response to these limitations, this paper introduces the Artificial General Intelligence Testbed (AGITB), a novel benchmark for validating foundational cognitive abilities in artificial intelligence systems.
-AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, thirteen of which are implemented as fully automatable and domain-agnostic tests.
+AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, with thirteen implemented as fully automated and domain-agnostic tests.
 Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour.
@@
 However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines.
-This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
+This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits, such as consciousness, self-awareness, and subjective experience, which remain elusive in artificial systems.
 Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have undergone significant shifts.
@@
 The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass1997.
-In alignment with this biologically grounded perspective, the AGITB departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level.
+In alignment with this biologically grounded perspective, the AGITB departs from symbolic, high-level evaluations and instead assesses intelligence at the lowest, signal-processing level.
 While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
@@
 A binary signal accurately represents the neural spike.
-Rooted in the view—advanced by Hawkins2004—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting.
-This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory spike trains, rather than disembodied symbols.
+Building on the view of Hawkins2004 that intelligence is fundamentally about extracting structure from data to enable prediction, AGITB operates at the level of signal-based forecasting.
+This approach aligns closely with the functioning of biological intelligence at the cortical level, which processes time-sensitive sensory spike trains rather than disembodied symbols.
 3 Artificial General Intelligence Testbed
@@
 A model must satisfy all requirements in order to claim success on the benchmark.
-The central premise is that the capacity to pass the AGITB constitutes a necessary condition for the existence of AGI.
-Non-AGI systems cannot meet AGITB’s requirements.
-Such a premise cannot be proven; it can only be rejected by any non-AGI system that meets the AGITB criteria.
-The fact that humans can pass the AGITB suggests that AGITB requires a solution beyond the current state-of-the-art AI.
-Unfortunately, without a clear definition of intelligence, we can only speculate that AGITB defines the necessary (but not sufficient) building blocks for the emergence of genuine, domain-independent intelligence.
-This assumption of necessity stems from the fact that, except for human brains, no existing artificial system has met its requirements.
-However, an AGITB-passing system is not automatically also a fully fledged AGI; after all, it is not required to express higher-level skills, like reasoning or natural language fluency.
-However, a failure to meet the AGITB’s requirements likely indicates a deficiency in foundational general cognitive capabilities.
-3.1 Components
-AGITB evaluates the user-supplied AGI model as a black box that predicts future inputs from the historical sequence of observed signals.
-Each input comprises ten bits, the specific semantics of which are immaterial; each bit may correspond to an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
+The central premise of AGITB is that it represents a necessary (though not sufficient) condition for the existence of artificial general intelligence.
+In the absence of a clear definition of intelligence, this premise cannot be formally proven, though it could be falsified by the existence of a non-AGI system that satisfies all AGITB criteria.
+The assumption of necessity follows from the fact that, apart from the human brain, no known artificial system has yet fulfilled its requirements, suggesting that a solution lies beyond the current state of the art in artificial intelligence.
+AGITB is not, in itself, a sufficient criterion for AGI; a system that passes all tests is not automatically a fully-fledged AGI, as it is not required to exhibit higher-level capabilities such as reasoning or natural language fluency.
+3.1 Architecture
+AGITB evaluates an AGI model as a black box that predicts the next input based on the historical sequence of observed signals.
+Each input consists of ten bits, the specific semantics of which are immaterial; each bit may represent an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
 Each input represents a snapshot of multiple parallel signals at a single time step (Figure 1a).
@@
 The interaction between spatial and temporal dimensions gives rise to structured patterns that are challenging for the model to adapt to.
-Figure 1: (a) Example of a 10-bit input $x$ with 4 bits set.
-(b) Example of an input sequence showing a temporal pattern of period four.
+Figure 1: (a) Example of a 10-bit input $x$ with four bits set.
+(b) Example of an input sequence.
 3.2 Operation
-The testbed presents the model with a stream of inputs over time.
-At each time step $t$, the model receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2.
+The testbed presents the AGI model with a stream of inputs over time.
+At each time step $t$, the model receives an input $x_{t}$ and is required to predict the subsequent input $x_{t+1}$, as shown in Figure 2.
 The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
 Figure 2: Iterative adaptation in discrete time.
-At the previous time step (t1t-1), the model issued a prediction $x_{t}^{*}$, with the second bit incorrect in this instance.
-After observing the realized input $x_{t}$, the model updates and subsequently produces a correct one-step-ahead prediction $x_{t+1}^{*}$ for the sequence shown in Figure 1b.
+At the previous time step (t1t-1), the model issued the prediction $x_{t}^{*}$.
+After observing the realised input $x_{t}$, it adapted its internal state in response to the error in the second bit and subsequently produced the one-step-ahead prediction $x_{t+1}^{*}$.
 3.2.1 Design
@@
 The reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
-Requirement 1 – Unbiased start
-Assertion: Models that have received no input are considered empty and thus equal.
-Assertion: An empty model predicts an empty input.
-AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs.
-Instead, they learn through interaction with their environment.
-Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics.
-While some reflexes may be genetically predetermined, these do not constitute genuine understanding.
-In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
-A model that predicts non-empty (non-zero) inputs prior to receiving data is considered biased, as it encodes unwarranted assumptions about the future.
-To ensure neutrality, models must initialise with empty predictions and remain unbiased until conditioned on actual input.
-Two model instances that have received no input should compare as equal, as neither has been influenced by prior experience.
-Here, "empty" refers to a lack of informational content, not structural capability.
-The model must possess an intrinsic organisational architecture capable of learning, even in the absence of prior data.
-The test operationalises, in empirical form, the philosophical challenge posed by classical symbolic systems: the emergence of meaning without prior grounding.
-Requirement 2 – Bias
-Assertion: Any model that has received input can no longer be regarded as unbiased.
+Requirement 1 — Unbiased start
+Assertion: A model starts unbiased.
+Assertion: An unbiased model produces an empty prediction.
+AGITB rests on the foundational assumption that general-purpose learning systems, such as the brain, do not begin with an innate understanding of external inputs but instead acquire meaning through interaction with their environment.
+Each system must construct semantic content from raw sensory data rather than rely on pre-encoded [^1] Although certain reflexes may be genetically specified, they do not constitute genuine understanding.
+As Hawkins2004 argue, low-level reflexes are not prerequisites for intelligence but rather evolutionary features of the subcortical “old brain”.
+knowledge.
+Accordingly, a model should originate in an unbiased initial state shaped exclusively by subsequent inputs, where “unbiased” denotes the absence of informational content rather than the lack of structural capacity for learning.
+A model that produces non-empty predictions prior to receiving any input is considered biased, as it encodes unwarranted assumptions about the future.
+To ensure neutrality, models must initialise in a completely blank state and remain so until conditioned on actual data.
+Although a spike-free prediction could be interpreted as a bias toward silence, this convention is justified by biological analogy: neurons do not activate in the absence of input.
+Two blank model instances should be regarded as equivalent, as neither has been influenced by prior experience.
+The unbiased start requirement serves as the foundation for other tests that operationalise the emergence of meaning in the absence of prior grounding.
+Requirement 2 — Bias
+Assertion: A model that has received input can no longer be regarded as unbiased.
 Every input modifies the model’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
-Requirement 3 – Determinism
-Assertion: If two cortices are equal, they must have received the same inputs.
+Requirement 3 — Determinism
+Assertion: If two models are equivalent, they must have received the same inputs.
 Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function.
-Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing.
-Similarly, in AGITB, two functionally identical models must have experienced identical input histories.
-Any difference in input must lead to a divergence in state, reinforcing the principle that behaviour and internal state are fully determined by input alone.
-The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave2016 in his discussion of free will.
-Determinism at the level of neural signal processing remains a necessary condition for the stable and reproducible functioning of the brain, even if its manifestations at the cognitive level are obscured by complexity and individual variation.
-Requirement 4 – Sensitivity
-Assertion: Two different cortices remain different, even if they experience long exposure to identical inputs.
-Cortical systems exhibit chaotic sensitivity to initial conditions.
-Small differences in early experiences or internal states can lead to divergent trajectories over time.
-This deterministic sensitivity, amplified through complex interactions, accounts for the illusion of unpredictability in decision-making.
-Requirement 5 – Time
-Assertion: Changing the input order results in a different model state.
-Because the model updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation.
-The ability to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
+Although minor stochastic effects may occur, they do not undermine the rule-governed nature of neural processing.
+By analogy, AGITB assumes that two functionally identical models must have experienced identical input histories: any difference in input necessarily produces a divergence in state.
+This principle reinforces the view that a model’s internal state is fully determined by its input history.
+Determinism at the level of neural signal processing is necessary for stable, reproducible brain function, whereas the apparent unpredictability of cognition stems from the system’s complexity rather than from genuine indeterminacy (Cave2016).
+Requirement 4 — Sensitivity
+Assertion: Two distinct models remain distinct, even after prolonged exposure to identical inputs.
+Cortical systems exhibit chaotic sensitivity to initial conditions: small variations in early experiences can lead to divergent trajectories over time.
+This deterministic sensitivity, amplified by complex internal interactions, underlies the apparent unpredictability of decision-making.
+Requirement 5 — Time
+Assertion: Changing the order of inputs results in a different model state.
+Since the model’s state depends on its cumulative history, the sequence of inputs critically shapes learning and adaptation.
+The capacity to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
 However, sensitivity to input order alone is not sufficient.
-Since temporal coding makes time an intrinsic property of neural data, Requirement 6 is designed to rule out models that perform equally well on temporally invariant data, as such models fail to exploit temporal information.
-Requirement 6 – Refractory period
-Assertion: The model must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
+Because temporal coding makes time an intrinsic feature of neural data, Requirement 6 is designed to exclude models that perform equally well on temporally invariant data, as such models fail to make functional use of temporal information.
+Requirement 6 — Refractory period
+Assertion: The model must be able to adapt to any minimal-period input sequence that respects biologically plausible refractory periods.
 Assertion: The model cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
-Biological intelligence depends on discrete spikes for signal transmission and learning.
-AGITB enforces refractory periods to reflect this constraint, capturing the biological principle that a neuron cannot fire again immediately after activation.
+Biological intelligence relies on discrete spikes for communication and learning.
+AGITB enforces refractory periods to reflect this constraint, embodying the biological principle that a neuron cannot fire again immediately after activation.
 While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing.
 This constraint helps preserve the temporal diversity needed for synaptic adaptation.
-Gerstner2002 showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning.
+Gerstner2002 showed that input sequences that lack sufficient variability, such as constant or overly repetitive spiking, fail to support effective learning.
 Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
-The refractory period requirement does not preclude the possibility of more general solutions.
-A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot simulate the absence of such dynamics.
-Requirement 7 – Temporal adaptability
+This requirement does not exclude the possibility of more general solutions.
+A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot its absence.
+Requirement 7 — Temporal adaptability
 Assertion: The model must be able to adapt to a temporal pattern with the user-specified period.
 Assertion: The model must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
-This test assesses the model’s ability to learn across different temporal scales.
-Unlike rigid temporal pattern-matching systems, intelligent models should be able to detect and respond to recurring structures, regardless of their exact periodicity.
-The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
-Requirement 8 – Stagnation
+This test requires the model to learn and track temporal structure across multiple timescales.
+Unlike rigid pattern-matching systems, an intelligent model should detect and predict a recurring structure regardless of its exact periodicity.
+Systems that can accommodate only a single, predetermined temporal scale therefore fail this requirement.
+Requirement 8 — Stagnation
 Assertion: There exists a limit beyond which the model can no longer adapt, even to patterns that would otherwise be learnable.
 Cognitive systems inevitably reach adaptation limits as their finite resources become saturated.
-This test assesses whether such limits emerge over time when the model is exposed to input sequences that would otherwise be learnable.
-Requirement 9 – Content sensitivity
+This test evaluates whether such a boundary arises when the model is exposed over time to input sequences that would otherwise be learnable.
+Requirement 9 — Content sensitivity
 Assertion: Adaptation time depends on the content of the input sequence.
-The structure and regularity of an input sequence influence the rate at which a model adapts, where adaptation time is defined as the number of iterations required to predict the entire temporal pattern accurately.
-Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
-Requirement 10 – Context sensitivity
+The structural complexity of an input sequence influences the rate at which a model adapts, where adaptation time is defined as the number of iterations required for the model to predict the entire temporal pattern accurately.
+Simple or highly regular sequences typically lead to rapid convergence, whereas irregular or noisy inputs demand longer exposure before the model can reliably capture and reproduce the underlying pattern.
+Requirement 10 — Context sensitivity
 Assertion: Adaptation time depends on the state of the model.
-Past learning influences how new information is integrated.
-A model with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganisation.
-Requirement 11 – Unobservability
-Assertion: Distinct cortices may exhibit the same observable behaviour.
-Identical external outputs can arise from different internal states.
-This many-to-one mapping underscores that observable behaviour alone cannot reveal the underlying structure or history of a model’s internal dynamics.
-Requirement 12 – Denoising
+The model’s current state reflects the cumulative influence of past inputs and therefore provides the context in which new information is processed.
+When subsequent inputs are consistent with structure already established through prior learning, adaptation proceeds quickly.
+Conversely, when new inputs conflict with this learned context, the model requires additional time to reorganise its state before accurate prediction becomes possible.
+Requirement 11 — Unobservability
+Assertion: Distinct models may exhibit the same observable behaviour.
+Identical external outputs may arise from distinct internal states.
+This many-to-one mapping highlights that observable behaviour alone cannot disclose the underlying structure or history of a model’s internal dynamics.
+Requirement 12 — Denoising
 Assertion: A model can recall a sequence despite perturbations.
-An intelligent model should be capable of efficiently memorising input patterns in a manner that enables the recall of previously observed sequences even when these are perturbed by noise.
+An intelligent model should be able to recall a previously observed sequence, even when the inputs are perturbed by noise.
 When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks.
-Average performance is used as the evaluation criterion, since random models may occasionally produce correct predictions by chance without reflecting genuine learning or structural understanding.
-Test 12 is designed as a deliberately conservative benchmark.
-The model is required to outperform random guessing in all 100 independent trials.
-This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it serves as a stringent robustness check.
-The design ensures that only large and systematic performance gains result in a pass.
-Modest or marginal improvements, while potentially real, are intentionally treated as failures, as the objective is to identify only clear and substantial advances in model capability.
+Average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
+Test 12 is intentionally designed as a conservative benchmark.
+The model must outperform random guessing in all 100 independent trials.
+This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it functions as a stringent robustness check.
+The design ensures that only large, systematic performance gains yield a passing result.
+Modest or marginal improvements, while potentially real, are intentionally regarded as failures, as the objective is to identify only clear and substantial advances in model capability.
 Requirement 13 — Generalization
@@
 After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs.
-Average performance is used as the evaluation criterion, since even a random model may occasionally produce correct predictions by chance without demonstrating genuine understanding or the acquisition of structured knowledge.
+As in the preceding requirement, average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
 The generalisation assessment follows the same conservative design as Test 12.
-By requiring success across all repetitions, the procedure establishes a stringent threshold whereby a single underperforming trial constitutes failure.
-Consequently, the test exhibits minimal tolerance for noise or marginal effects.
-This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models demonstrating a clear and consistent advantage achieve a passing result.
-Requirement 14 – Bounded Prediction Latency
+By requiring success across all repetitions, the procedure imposes a stringent threshold in which a single underperforming trial results in failure.
+As a consequence, the test tolerates virtually no noise or marginal effects.
+This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models exhibiting a clear and consistent advantage obtain a passing result.
+Requirement 14 — Bounded Prediction Latency
 A model must provide low-latency, signal-level predictions to ensure functional viability.
-Biological brains achieve approximately constant reaction times through extensive parallel processing, such that per-input latency remains bounded and stable under expected workloads.
-No universal automated test executed on serial hardware can feasibly verify this property across all possible model implementations.
-Consequently, model developers are responsible for ensuring that their architectures permit efficient parallel execution with bounded latency.
-In this regard, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
+Biological brains achieve approximately constant reaction times through extensive parallel processing, keeping per-input latency bounded and stable under expected workloads.
+No universal automated test running on serial hardware can feasibly verify this property across all possible model implementations.
+Accordingly, it is the responsibility of model developers to ensure that their architectures support efficient parallel execution with bounded latency.
+In this respect, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
 3.4 Search space
 To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory.
-In AGITB, each task typically involves predicting a random temporal sequence of seven binary inputs, each comprising ten bits.
-This defines a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
-The refractory period is a biologically inspired constraint that prohibits any neuron (bit) from firing in consecutive time steps.
-This restriction substantially reduces the number of valid input sequences.
-There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits, assuming that a $1$ never carries over to the next time step, where $F_{i}$ is the $i$-th Fibonacci number with $F_{0}=0$.
+In AGITB, tasks typically involve predicting a random temporal sequence of seven (or more) binary inputs, each comprising ten bits.
+This yields a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
+AGITB incorporates a biologically inspired refractory period, which prohibits any neuron (bit) from firing in consecutive time steps.
+This restriction substantially reduces the number of valid sequences.
+There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits under the condition that a $1$ never carries over to the next time step, where $F_{i}$ denotes the $i$-th Fibonacci number with $F_{0}=0$.
 In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence.
-The number of distinct cyclic temporal input sequences with the refractory constraint is $|S"|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
-The use of length-seven sequences with ten-bit inputs is considered sufficient to detect non-AGI behaviour while maintaining computational efficiency.
-Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing execution time.
-The current configuration ensures that each test remains computationally feasible and diagnostically informative.
+The number of distinct cyclic temporal sequences respecting the refractory constraint is $|S^{\prime\prime}|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
+The choice of seven-step sequences with ten-bit inputs is sufficient to detect non-AGI behaviour while maintaining computational efficiency.
+Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing runtime.
+The current configuration therefore ensures that each test remains both computationally feasible and diagnostically informative.
 Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure.
-However, the sheer size of the search space renders explicit optimisation —or "teaching to the test"—computationally infeasible.
-Given that real-world sensory inputs may encompass tens of thousands of bits, a genuine AGI system must therefore employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from input data.
+However, the sheer size of the search space makes any form of explicit teaching-to-the-test computationally infeasible.
+Given that real-world sensory inputs may ultimately encompass tens of thousands of bits, a genuine AGI system must employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from high-dimensional data.
+3.5 Performance
 3.6 Performance
 Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself.
-Similar to the Turing Test—which Harnad1992 described as an empirical criterion rather than a philosophical sleight of hand—AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence.
-The ultimate objective remains the development of AGI, not merely success on the benchmark.
-AGITB can yield meaningful insights only when developers adhere strictly to its core requirements.
-Misinterpretations of fundamental elements, such as the "empty" initial state, may lead to erroneous conclusions and hinder genuine progress toward AGI.
+Similar to the Turing Test, which Harnad1992 characterised as an empirical criterion rather than a philosophical manoeuvre, AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence.
+The ultimate objective remains the development of AGI, not merely achieving success on the benchmark.
+AGITB yields meaningful insights only when developers adhere strictly to its core requirements.
+Misinterpretations of fundamental elements, such as the notion of an “empty” initial state, may lead to erroneous conclusions and impede genuine progress toward AGI.
 Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models.
-However, before benchmarking artificial systems, it is necessary to consider the baseline performance of human cognition.
+Before benchmarking artificial systems, however, it is necessary to consider the baseline performance of human cognition.
 3.6.1 Human Performance
 The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense.
-Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with cognitive expectations, it can be assumed that humans naturally satisfy most requirements.
+Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with basic cognitive competencies, it is reasonable to assume that humans naturally satisfy most requirements.
 Demand 1 (Unbiased start), however, warrants further discussion.
-Owing to prior experience and cognitive bias, an adult human may appear to "fail" the unbiased start prerequisite, as the cortex is no longer in an unconditioned state and may generate non-empty predictions.
-However, AGITB specifies that the unbiased state must precede the first input, a condition met in the fetal cortex.
-At that stage, the human cortex lacks synaptic organisation and, upon receiving its first sensory input, would satisfy the criterion of true neutrality by predicting no spikes.
-Several of the more complex AGITB tests have cognitive-level analogues that can be observed through conscious reasoning and introspection.
-Temporal flexibility (Requirement 7), for example, is straightforward for humans, who can easily recognise temporal patterns of varying durations.
-Stagnation (Requirement 8) reflects the physical limitation of the human cortex to store and maintain knowledge.
-The corresponding test is analogous to the onset of cognitive saturation or early dementia, when recent experiences are forgotten.
-Because humans learn different topics at varying speeds, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is self-evident.
-The unobservability test (Requirement 11) captures the fact that internal human mental states cannot be directly inspected.
-The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when faced with new or distorted inputs.
-Finally, bounded prediction latency (Requirement 14) is a defining property of the human brain.
+Owing to prior experience and cognitive bias, an adult human’s cortex may appear to “fail” the unbiased start prerequisite, as it is no longer in an unconditioned state and may generate non-empty predictions.
+AGITB, however, requires that the unbiased state precede the first input—a condition satisfied only in the fetal cortex.
+At that stage, the cortex lacks synaptic organisation and, upon receiving its first sensory input, meets the criterion of true neutrality.
+Several of the more complex AGITB tests have cognitive-level analogues that can be observed through reasoning and introspection.
+Temporal flexibility (Requirement 7), for example, poses no difficulty for humans, who readily recognise temporal patterns of varying durations.
+Stagnation (Requirement 8) reflects the finite capacity of the human cortex to store and maintain knowledge; its behavioural analogue resembles the onset of cognitive saturation or early dementia, in which recent experiences are lost.
+Because humans learn different types of content at varying rates, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is immediately evident.
+The unobservability test (Requirement 11) formalises the fact that internal human mental states cannot be directly inspected.
+The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when confronted with new or distorted inputs.
+Finally, bounded prediction latency (Requirement 14) is a well-established property of the human brain, which maintains approximately constant reaction times through extensive parallel processing.
 3.6.2 Classical Symbolic Programs
 In principle, two complementary designs for AGI can be conceived.
-The first encodes, either explicitly or implicitly, certain expectations about the external world within the system itself.
-The second is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
-The former category encompasses most AI and putative AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990.
-Although such systems may exhibit behaviour that appears intelligent, they lack grounded understanding and therefore cannot be regarded as genuine instances of AGI.
-More specifically, systems that embed expectations inevitably reflect the programmer’s assumptions about the meaning of the signals from which they learn.
-In the case of standard symbolic programming, the program itself constitutes prior knowledge—its rules and representations presuppose an interpretation of the symbols being manipulated.
-The existence of such a program directly violates AGITB’s first test, which prohibits the inclusion of any external knowledge.
-Namely, the AGI program itself introduces the symbol-grounding problem it is meant to overcome.
-AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary inputs.
+One embeds, either explicitly or implicitly, expectations about the external world within the system.
+The other is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
+The former category includes most AI and purported AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990.
+Although such systems may display behaviour that appears intelligent, their interpretations of symbols depend on programmer-supplied conventions rather than grounded understanding, and they therefore cannot qualify as genuine AGI.
+More specifically, systems that embed expectations inevitably incorporate the designer’s assumptions about the meaning of the signals they process.
+In classical symbolic architectures, the program itself constitutes prior knowledge: its rules and representations presuppose interpretations of the symbols being manipulated.
+The very existence of such a program violates AGITB’s first test, which prohibits external knowledge of any kind.
+In effect, the AGI program smuggles in the symbol-grounding problem it is meant to avoid.
+AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary signals.
 Only an expectation-free system could, in principle, satisfy this condition.
-To date, however, no such program is known.
-This suggests that a true AGI may not explicitly encode the operations of intelligence itself, but rather the dynamics of the substrate from which intelligence can emerge—a view consistent with the "Brain Simulator Reply" proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980).
+This perspective suggests that a genuine AGI may not explicitly encode the operations of intelligence, but rather the dynamics of a substrate from which intelligence can emerge—a view aligned with the “Brain Simulator Reply” proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980).
+To date, however, no such system is known.
 3.6.3 Artificial Neural Network Performance
-A similar dichotomy applies to connectionist architectures.
-Artificial neural networks can, in principle, be constructed either with built-in expectations—through pretraining and architectural priors—or as expectation-free systems that begin from a neutral state.
-However, AGITB’s requirement that a model start in a completely unbiased, knowledge-free condition stands in fundamental tension with the prevailing paradigm of modern deep learning.
-Contemporary neural models typically rely on pretraining, whereby network weights are adjusted in advance through exposure to structured or labelled data.
-Only an untrained network can be regarded as unbiased.
-However, an untrained network with all weights initialised to zero exhibits no learning dynamics: the neurons are not halted but remain dormant, producing no output.
-This exposes a fundamental limitation of current ANN architectures—they lack an intrinsic mechanism for initiating unbiased, autonomous learning from scratch in arbitrary and unfamiliar environments.
-Furthermore, by translating symbols into numerical vectors, ANNs effectively transform the symbol grounding problem into a number grounding problem.
-While these numerical representations capture relational regularities among symbols, they also introduce spurious associations that are not anchored in real-world semantics, giving rise to the phenomenon commonly referred to as hallucination.
-Pretraining on large datasets establishes correlations rather than genuine understanding.
-AGITB rejects this approach in favour of internal learning algorithms capable of initiating adaptation autonomously.
-However, no such mechanism has yet been demonstrated.
+A similar dichotomy applies to connectionist architectures: ANNs can, in principle, be constructed either with built-in expectations (introduced through pretraining regimes or architectural priors) or as expectation-free systems that begin from a neutral initial state.
+AGITB’s requirement that models begin in a completely unbiased, knowledge-free state stands in fundamental tension with the dominant paradigm of modern deep learning.
+Contemporary neural models typically rely on extensive pretraining, during which network weights are shaped by prior exposure to structured or labelled data.
+Moreover, by mapping symbolic inputs to numerical vectors, standard ANNs effectively shift the symbol-grounding problem into a number-grounding problem.
+Although these vector representations capture relational regularities within the training data, they also introduce spurious associations not anchored in real-world semantics, leading to the phenomenon commonly described as hallucination.
+Only an expectation-free network can be regarded as unbiased.
+Yet a network without any initialised weights exhibits no effective learning dynamics: its units are not halted, but remain dormant and produce no informative output.
+This highlights a fundamental limitation of current ANN architectures: they do not initiate learning autonomously but instead depend on an external training procedure to drive adaptation.
+AGITB, by contrast, requires a blank system capable of initiating adaptation autonomously in an unfamiliar environment.
+To date, no such mechanism has been demonstrated in artificial neural networks.
 3.6.4 Large Language Model Performance
-Because a large language model (LLM) is a type of artificial neural network (ANN), it fails the unbiased start test (Requirement 1).
-This failure demonstrates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained.
-Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—such evaluations remain moot until the first requirement is satisfied.
+Because a large language model is a type of artificial neural network, it fails the unbiased start test (Requirement 1).
+LLMs cannot learn entirely from scratch, as their behaviour is shaped in advance by the linguistic patterns and data on which they were pretrained.
+Although an LLM’s internal state is readily accessible and, unlike that of a human, can be inspected or compared across instances, any such evaluations remain moot until the first requirement is satisfied.
 Two possibilities warrant consideration.
-The first concerns whether LLMs, while not qualifying as AGI, might nevertheless simulate one.
-For example, an LLM could in principle be prompted to imitate a model receiving binary inputs as specified by AGITB; a structured prompt of this type is provided in the Appendix.
-However, before evaluating such simulations, it is necessary to revisit the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs.
-This condition is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window.
-Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to lose or attenuate the long-term dependencies required by the test.
-The second possibility is whether an LLM can autonomously generate an architecture capable of satisfying AGITB, given the reference implementation.
-In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purporting to pass the tests, none advanced beyond the adaptability test (Requirement 7).
-To summarise, LLMs do not perform genuine learning solely from prompts, nor can they develop the grounded, context-dependent understanding of the world characteristic of human cognition.
-These limitations extend to large reasoning models (LRMs), which share the same underlying architectural principles.
+The first is whether LLMs, while not constituting AGI, might nevertheless simulate one.
+For example, an LLM could be prompted to imitate a model receiving binary inputs as specified by AGITB.
+However, before examining such simulations, it is necessary to consider the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs.
+This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window.
+Once this window is exceeded, earlier tokens are discarded or compressed, weakening or eliminating the long-term dependencies required by the test.
+The second possibility is whether an LLM can autonomously generate a solution to AGITB when provided with the reference implementation.
+In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purported to satisfy the requirements, none progressed beyond the adaptability test (Requirement 7).
+In summary, LLMs do not perform genuine learning solely from prompts, nor can they acquire the grounded, context-dependent understanding characteristic of human cognition.
+These limitations extend to large reasoning models, which share the same underlying architectural foundations.
 3.7 Remarks
-AGITB evaluates the model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs.
-Random input sequences with arbitrary internal correlations are used to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge.
-By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
-The low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador2023, where cognitive understanding emerges from the integration of continuous sensory streams.
-The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT.
-Hand-engineered, task-specific systems have historically struggled to scale toward true generalisation, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks.
-However, state-of-the-art AI networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input.
-This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB?
+AGITB evaluates a model’s predictive capabilities after exposure to temporal sequences of both structured and random inputs.
+Random input sequences with arbitrary internal correlations are employed to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream itself rather than from prior knowledge.
+By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals; the random inputs need not resemble real-world sensory data.
+The low-level, binary operational framework makes AGITB particularly well suited for evaluating NeuroAI models that aim to satisfy principles of the embodied Turing Test, as proposed by Zador2023, wherein cognitive understanding emerges from the integration of continuous sensory streams.
+The progression from raw signal prediction to higher-level abstraction mirrors the broader trajectory of AI, from early perceptrons to large-scale models such as GPT.
+Hand-engineered, task-specific systems have historically struggled to scale toward genuine generalisation, underscoring the need for adaptive, learning-based architectures such as artificial neural networks.
+However, state-of-the-art neural systems rely heavily on pretraining with structured or symbolic data, a process that inevitably introduces biases and bypasses the grounding of meaning in raw sensory input.
+This raises a central and unresolved question: how can such models initiate learning directly from unstructured input, as required to satisfy AGITB?
 3.8 Cheating the Benchmark
-Since AGITB tests are simple to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific solutions and make the model deploy them selectively based on the task at hand.
-While it is theoretically possible to infer the model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research.
-Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the programming interface—these measures would not only deter cheating but also make tests less intelligible to human developers.
-This would undermine the testbed’s core objectives: transparency, readability, and interpretability.
-For this reason, the AGITB reference implementation remains deliberately readable and accessible.
-To date, no non-AGI system has demonstrated the level of performance that AGITB demands.
-Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
+Because AGITB’s tests are individually simple to solve, one might imagine circumventing the benchmark by engineering task-specific solutions and having the model selectively deploy them depending on the detected test scenario.
+In principle, the task being administered could be inferred by monitoring the number of instantiated models and the sequence of invoked methods.
+However, such an approach would amount to subverting the benchmark rather than advancing AGI research.
+Although AGITB could be hardened against this form of cheating (by, for example, shuffling tests or redesigning the programming interface) these measures would reduce the transparency and interpretability of the testbed and thereby hinder its intended use by human developers.
+For this reason, the AGITB reference implementation is kept deliberately readable and accessible.
+To date, no artificial system has demonstrated the level of performance that AGITB requires.
+Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing a more obfuscated version of the testbed.
 4 Competing Benchmarks
-Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation.
-A related effort is NeuroBench, which is designed to support the development of neuromorphic approaches to AGI.
-Both ARC and NeuroBench employ a variety of correctness and complexity metrics to differentiate among non-AGI systems, as they are primarily intended to distinguish weaker from stronger narrow models.
-In contrast, AGITB is designed to identify only true positive instances of AGI.
+Among existing benchmark tasks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation.
+A related effort is NeuroBench, which is designed to support the systematic evaluation of neuromorphic and other biologically inspired architectures.
+Both ARC and NeuroBench rely on a variety of correctness and complexity metrics to compare non-AGI models; their primary purpose is to distinguish weaker from stronger narrow systems.
+In contrast, AGITB is constructed to assess whether a model satisfies a set of necessary conditions associated with artificial general intelligence, rather than to rank models within a specific performance spectrum.
 4.1 ARC
-ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recolouring, rearranging, or modifying spatial patterns in a spatial grid.
-However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference.
-These priors are left undefined, placing an ambiguous burden on the model architecture.
-In contrast, AGITB adopts a fundamentally different perspective: it treats the model under evaluation as a blank slate that must develop structure and function exclusively through interaction with temporally structured input.
-Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical "old brain," as described by (Hawkins2004, p.
-66).
-While ARC presumes temporal reasoning, it does not explicitly test it, as it uses only two images to demonstrate a transformation.
-In contrast, AGITB evaluates cognition as a dynamic, unfolding process over time.
-A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure.
-For instance, to learn about objects and to recognise the object transformation, such as "move left," a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything.
-Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
-In summary, ARC requires a high-level intelligence, whereas AGITB requires the 14 low-level principles.
-ARC evaluates performance based on behaviours grounded in human-like cognitive priors, while AGITB determines a system’s ability to support the emergence of these priors through exposure to raw input.
-Benchmarks such as ARC and the Turing test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria.
-In contrast, AGITB conceptualises intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions.
+ARC presents visual reasoning tasks in which a model must infer novel transformations (such as recolouring, rearranging, or modifying spatial patterns) from a sequence of two input–output examples defined on discrete spatial grids.
+However, ARC implicitly assumes the presence of high-level cognitive priors, including object permanence, spatial reasoning, numerical abstraction, and causal inference.
+These priors are not formally specified, placing an ambiguous and open-ended burden on the model designer.
+In contrast, AGITB adopts a fundamentally different stance: it treats the system under evaluation as a blank slate that must acquire structure and function exclusively through interaction with temporally structured input.
+Although ARC presumes temporal reasoning, it does not explicitly test it, since each task presents only two images to illustrate a transformation.
+AGITB, by contrast, evaluates cognition as a dynamic process unfolding over time.
+A model can acquire knowledge and predictive capability only through continuous exposure to temporally structured data, not from disconnected before–after snapshots that lack the temporal continuity needed to infer causal relationships.
+For example, to recognise an object moving left, a model in AGITB must observe multiple intermediate states across time; the final image alone is insufficient to infer the transformation.
+Temporal structure, rather than static pattern comparison, provides the substrate from which invariants and causal relations can be learned.
+In summary, ARC evaluates high-level intelligence grounded in human cognitive priors, whereas AGITB evaluates adherence to fourteen low-level computational principles intended to support the emergence of such priors.
+ARC and the Turing Test both frame intelligence through an anthropocentric lens, embedding assumptions drawn from human cognition.
+AGITB instead conceptualises intelligence as a universal capacity for learning that does not rely on innate symbolic structures or species-specific expectations.
 4.2 NeuroBench
 NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics.
-It is particularly specialised for neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability.
-Within its algorithm track, the framework evaluates model performance on several ongoing challenges in general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
+It is particularly oriented toward neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability.
+Within its algorithm track, the framework evaluates models on several challenges relevant to general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
 The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states.
-For this purpose, NeuroBench employs a synthetic one-dimensional Mackey–Glass time series—a dataset specifically designed for architectures with limited input/output capacity.
-However, several issues limit NeuroBench’s suitability as a general AGI benchmark.
-First, the Mackey–Glass data are numerical; the meaning of a number, much like that of a symbol, is defined by human interpretation.
-The use of such representations necessarily reintroduces the symbol-grounding problem.
-Second, the threshold corresponding to AGI-level performance remains undefined.
-Although the symmetric mean absolute percentage error (sMAPE) is an established metric in forecasting and NeuroBench provides a baseline score, it does not specify what level of accuracy constitutes general intelligence.
-After all, humans perform poorly at detecting anomalies or making long-term predictions of the Mackey–Glass signal (Thill2020).
-Third, although long-term prediction is not problematic in itself, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to recognise and correct its own errors.
-This design is incompatible with online learning, where an AGI should continuously update its internal state upon detecting discrepancies between its predictions and the actual outcomes.
-NeuroBench, by contrast, implicitly assumes that an AGI would behave as a purely mechanical predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, or genuine agency.
+NeuroBench employs a synthetic one-dimensional Mackey-Glass time series for this task, a dataset designed for architectures with limited input/output capacity.
+However, several issues limit the usability of chaotic function prediction (CFP) as a general AGI benchmark task.
+First, the Mackey-Glass data are numerical, and NeuroBench does not prescribe the encoding scheme.
+An inappropriate encoding can distort the temporal and causal structure of the observed signals, such that a numeric value—much like a symbol—derives its meaning from human interpretation rather than from the model’s own grounded understanding.
+This effectively reintroduces the symbol-grounding problem in a numerical form.
+Second, the threshold for AGI-level performance is not clearly defined.
+Although symmetric mean absolute percentage error (sMAPE) is a standard forecasting metric, NeuroBench does not specify what performance level would correspond to general intelligence.
+Notably, humans themselves perform poorly at anomaly detection and long-horizon prediction of the Mackey-Glass signal (Thill2020).
+Third, although long-term prediction is not inherently problematic, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to detect and correct its own errors.
+This design is incompatible with online learning, in which an AGI should continuously update its internal state upon observing discrepancies between prediction and outcome.
+NeuroBench, by contrast, emphasises offline learning and assumes that an AGI would behave as a purely mechanistic predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, and genuine agency.
 4.3 Direct comparison
-Table 1 summarises the principal distinctions among the three benchmarks.
-ARC and NeuroBench each presuppose the presence of high-level cognitive components—visual object recognition, manipulation and reasoning in the case of ARC, and comprehension and reasoning abilities in the case of NeuroBench.
-AGITB, by contrast, specifies elementary requirements that can be evaluated directly.
+Table 1 highlights the key differences among the tasks used in the three benchmarks.
+Whereas ARC and NeuroBench presuppose or require models to exhibit high-level cognitive capacities (such as object recognition, spatial manipulation, and various forms of reasoning), AGITB instead focuses on minimal, precisely defined requirements that can be evaluated directly at the signal-processing level.
 Property ARC CFP AGITB
 Interface modality Visual Numeric Binary
-Target reasoning type Human Universal Universal
-Assumed cognitive priors Yes No No
-Automation Limited Yes Yes
-Teaching-to-the-test Yes No No
-Intelligence level High Medium Low
+AGI type Human Universal Universal
+Cognitive priors Yes No No
+Abstraction level High Medium Low
+Formal AGI requirements No (qualitative) Partial Yes (explicit, testable)
 Task preparation Manual Automatic Automatic
 Symbol Grounding Problem Yes Yes No
-Spatial dimension $200\times 100$ 1 10
-Temporal dimension Yes Yes Yes
-Table 1: A list of core properties of the ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
+Input dimensionality $200\times 100$ (4 colours) 1 number 10 bits
+Temporal sequence length 2 750+ 7+
+Table 1: Core properties of ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
 5 Conclusion
-Although AGITB is proposed as a general benchmark for artificial general intelligence, it is evidently grounded in empirical knowledge of neural function in the human brain.
-This grounding is justified by the fact that we have evidence that neuron-based systems—such as biological brains and artificial neural networks—are capable of producing general intelligence.
-In contrast, we have no demonstrated examples of AGI arising from fundamentally different architectures.
-Therefore, seeking more abstract or universal criteria may be premature, at least until the principles underlying natural, neuron-based intelligence are more fully understood.
-Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex.
+Although AGITB is proposed as a general testbed and benchmark for artificial general intelligence, it is nonetheless grounded in empirical knowledge of neural function in the human brain.
+This grounding is motivated by the fact that neuron-based systems—most notably biological brains and, to a more limited extent, artificial neural networks—are the only systems known to support a broad range of flexible cognitive abilities.
+By contrast, there is currently no evidence that fundamentally different computational architectures can give rise to comparable forms of general intelligence.
+Pursuing more abstract or architecture-agnostic criteria may therefore be premature until the principles underlying natural, neuron-based intelligence are better understood.
+Unlike conventional benchmarks that focus on high-level task performance, such as question answering or language translation, AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex.
 It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence.
-The proposed testbed introduces a systematic framework comprising twelve fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
-AGITB requires models to begin from an unbiased initial state and to develop functionality solely through exposure to structured or unstructured input.
-This design reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity.
-In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals.
-This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalisation over time.
-AGITB is solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems.
-This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence.
-While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence.
-As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
-// TODO While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture that can satisfy all required tasks across the full suite within a cohesive and unified framework.
-As a recently proposed benchmark, AGITB has not yet been passed by any non-AGI system, and it remains an open question whether such a solution is even theoretically possible.
+The proposed testbed introduces a systematic framework comprising fourteen fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
+Unlike conventional benchmarks that target high-level task performance, such as question answering or language translation, AGITB evaluates whether a system exhibits behaviours associated with core operational principles of the biological cortex.
+Its focus is on low-level, biologically grounded computational properties that are believed to underlie the emergence of general intelligence.
+The testbed comprises fourteen tightly interdependent tests, each simple in isolation but collectively requiring the kind of adaptive, general-purpose learning expected of an AGI.
+AGITB requires models to begin in an unbiased initial state and to acquire all functionality solely through exposure to structured or random input.
+This mirrors neuroscientific evidence that cortical learning is fundamentally input driven: neural circuits develop through experience, not through pre-encoded semantics.
+In biological systems, high-level cognition arises not from symbolic manipulation but from the continual adaptive prediction of low-level sensory signals.
+Such prediction is more than pattern matching; it supports the progressive construction of signal-grounded knowledge from which abstraction and generalisation can emerge.
+AGITB is solvable by humans yet remains unsolved by classical algorithms and current state-of-the-art AI systems.
+This persistent performance gap provides empirical support for the claim that AGITB targets capabilities characteristic of general rather than narrow intelligence.
+Although the absence of an artificial solution does not constitute a formal proof of adequacy, the fact that humans succeed where machines do not indicates that the benchmark captures functionally relevant aspects of general intelligence.
+In this sense, AGITB serves as a discriminative test and a principled tool for steering the development of systems capable of genuinely general, adaptive learning.
 Acknowledgments and Disclosure of Funding